### v1

# Robust Few-shot Transfer Learning for Knowledge Base Question Answering with Unanswerable Questions

Abstract
Real-world KBQA applications require models that are (1) robust – e.g., can differentiate between answerable and unanswerable questions, and (2) low-resource – do not require large training data. Towards this goal, we propose the novel task of few-shot transfer for KBQA with unanswerable questions. We present FUn-FuSIC that extends the state-of-the-art (SoTA) few-shot transfer model for answerable-only KBQA to handle unanswerability. It iteratively prompts an LLM to generate logical forms for the question by providing feedback using a diverse suite of syntactic, semantic and execution guided checks, and adapts self-consistency to assess confidence of the LLM to decide answerability. Experiments over newly constructed datasets show that FUn-FuSIC outperforms suitable adaptations of the SoTA model for KBQA with unanswerability, and the SoTA model for answerable-only few-shot-transfer KBQA.
1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes a Knowledge Base (KB) and a natural language question as input and requires as output a logical form (or program) that produces an answer upon execution over the KB. KBQA has important real-world applications, which require KBQA systems to be low-resource (requiring only a few task-specific labeled examples) and robust. One important aspect of robustness is the ability to identify unanswerable questions, which arise naturally in real applications Patidar et al. (2023). Traditional supervised models Saxena et al. (2022); Zhang et al. (2022); Mitra et al. (2022); Wang et al. (2022); Das et al. (2022); Ye et al. (2022); Chen et al. (2021); Das et al. (2021); Shu et al. (2022); Gu et al. (2023) and even recent LLM-based few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) are found wanting in both aspects.
Recent work has addressed low-resource KBQA in the form of few-shot transfer learning Patidar et al. (2024) – the idea is to exploit existing KBQA datasets (source) and transfer them few-shot to a target setting. Their FuSIC-KBQA architecture uses supervised KB retrieval followed by generation and refinement of logical forms using LLM few-shot in-context learning.
A parallel direction addresses unanswerable questions for KBQA Patidar et al. (2023). Unanswerability arises due to deficiencies in the schema or the data in the KB.  Faldu et al. (2024) proposes RetinaQA – a supervised architecture (with high training data needs) that is robust to both answerable and unanswerable questions. However, no existing architecture simultaneously addresses the low-resource and robustness requirements.
In response, we propose the task of few-shot transfer learning for KBQA in the unaswerability setting. Specifically, the target KBQA task offers only a few labeled examples, while the source KBQA task has thousands of labeled examples containing answerable questions only.
Building a model for this setting could involve extending FuSIC-KBQA for unanswerability or RetinaQA for few-shot transfer. Since RetinaQA’s design requires large volumes of labeled training data in the target and it fares poorly in few-shot setting, we build our model on top of FuSIC-KBQA. A natural vanilla extension is to modify the LLM prompt for unanswerability and including different types of unanswerable questions in the few shots, as proposed in the broader unanswerable QA literature Slobodkin et al. (2023). Unfortunately, we find that this largely fails for unanswerable questions that do not have a valid logical form – the simple modification regularly labels these as answerable and outputs incorrect logical forms. A second issue is that FuSIC-KBQA prompts LLM to refine the logical form if its execution leads to an empty answer – but this assumes that all questions are answerable and naturally breaks down in the face of unanswerability.
To handle this, we develop FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA). FUn-FuSIC takes the vanilla approach further by detecting when the LLM is very uncertain about the logical form, which is suggestive of the question being unanswerable. For this, it employs logical form refinement by providing LLM repeated feedback about potential errors, including syntactic, semantic and execution-related error checks, and a novel strategy of back-translating the logical form to check equivalence with the original question. To assess the model’s lack of certainty about the logical form, FUn-FuSIC adapts self-consistency Wang et al. (2023) – in the absence of enough certainty, schema-level unanswerability is inferred. On the other hand, when a sufficient number of logical forms agree on the answer (empty or non-empty), the LLM is finally prompted to choose the best one, along with inferring answerability (for non-empty answer) or data-related unanswerability (for empty answer).
For evaluation, we first create two new datasets for few-shot transfer for KBQA with unanswerability. Using these, we show that, in aggregate, FUn-FuSIC comprehensively outperforms RetinaQA, when adapted for few-shot transfer, as well as the vanilla extension to FuSIC-KBQA.
We further demonstrate that the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks, holds promise for KBQA in general. Using experiments over existing few-shot transfer datasets for answerable KBQA, we find that the restriction of FUn-FuSIC for the answerable setting outperforms FuSIC-KBQA, which is the SoTA KBQA model in this setting.
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative feedback for error checks and self-consistency adapted for unaswerability. (c) We create new datasets for the proposed task which we make public  [^1] https://anonymous.4open.science/r/FUn-FuSIC-1704. (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting. (e) We also show that in the standard answerable setting, FUn-FuSIC outperforms the corresponding SoTA model.
2 Related Work
There is a lot of work on in-domain KBQA using supervised models  Ye et al. (2022); Chen et al. (2021); Shu et al. (2022); Gu et al. (2023) as well as LLM few-shot approaches Li et al. (2023); Nie et al. (2024); Shu and Yu (2024). These require significant volumes of labeled data to train or select the most appropriate few shots.
Also in the in-domain setting, unanswerability has been studied for KBQA Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) created the GrailQAbility dataset with different categories of unanswerability, and showed that answerable-only models, when adapted superficially for unaswerability, are inadequate in different ways. RetinaQA Faldu et al. (2024) is the SoTA model for unaswerability. It learns to discriminate between valid logical forms and those rendered invalid due to schema-level incompleteness, and constructs logical forms top-down using sketch-filling to handle data-level incompleteness. However, this also requires large volumes of training data.
In the transfer setting for KBQA, one formulation makes use of large volumes of unlabeled training data in the target Cao et al. (2022); Ravishankar et al. (2022). More recently, FuSIC-KBQA has addressed the few-shot transfer setting with only a small volume of labeled training examples in the target Patidar et al. (2024). This uses a retrieve-then-generate framework using one or more supervised retrievers and LLM few-shot in-context learning based generators. The LLM generation also goes through multiple iterations of execution-guided feedback until a non-empty answer is obtained. However, the transfer setting has only considered answerable questions.
LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al. (2023), but without feedback for iterative generation. Other approaches  Shinn et al. (2023); Chen et al. (2023b) use execution based refinement for programs but do not involve unanswerability.
In the prompting literature, self-consistency marginalizes over multiple reasoning paths to arrive at a consensus answer Wang et al. (2023); Chen et al. (2023a). This works only when significant probability mass accumulates for at least one answer, which is selected as the winner. This breaks down under unaswerability, when reasoning paths to all non-empty answers have low probability.
3 Problem and Proposed Approach
Figure 1: Illustration of FUn-FuSIC for a specific question when executed over three different KBs. The question is answerable for KB3, but unanswerable for KB1 and KB2, due to schema incompleteness and data incompleteness respectively. The iterations as shown as i=1,2,3,4, The consistency block is below these. Colored boxes show different checks. Additions to the set of candidate logical forms are shown in the gray columns.
We first briefly cover background on KBQA, unaswerability and few-shot KBQA. More details are in the Appendix (Sec. A.1). We then define the few-shot setting with unanswerability and our proposed approach.
3.1 Few-shot Transfer Learning for KBQA with Unanswerability
A Knowledge Base $G$ consists of a schema with data stored under it. The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. We are given a target Knowledge Base $G^{t}$ and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns an answer set $a^{t}$. We are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms. A related source domain has a source knowledge base $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples. The source and target KBs and the distributions of questions and logical forms can be very different.
In general, a KB may have both answerable and unanswerable questions Patidar et al. (2023). A question $q$ is answerable for a KB $G$ if it has a corresponding logical form which when executed over $G$ returns a non-empty answer $A$. In contrast, a question is unanswerable for $G$, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form for $G$, but which on executing on $G$ returns an empty answer (data-level unanswerability). Schema-level unaswerability arises due to missing types (classes) and relations. On the other hand, data-level unaswerability arises due to missing entities and facts. Absence of any mentioned question entity, however, invalidates the logical form. More details are in the Appendix (Sec.  A.1.3).
In KBQA with unanswerability, given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
We now propose the problem of few-shot transfer learning for KBQA with unanswerability. Here, the target KB $G^{t}$ is incomplete, and target test questions $q^{t}$ may be answerable or unanswerable due to missing schema or data. The few shot examples $D^{t}$ contain both answerable and unanswerable questions of different types. Since almost all large KBQA datasets contain only answerable questions, we assume that the source training data $D^{s}$ contains only answerable questions. So, in addition to the differences between source and target KBQA tasks in the existing transfer learning setting, the unanswerability gap also needs to be bridged.
3.2 Proposed Approach: FUn-FuSIC
Our proposed architecture FUn-FuSIC builds upon and adapts FuSIC-KBQA for unaswerability. We preserve its retrieve-then-generate framework where the first stage is KB retrieval with one or more supervised retrievers, which are source-trained and then target fine-tuned. The retrieval is LLM reranked to increase target relevance. This forms the input to the generative stage, which prompts an LLM to generate the logical form using the retrieval, either in few-shot or zero-shot fashion. More details of these stages are in the Appendix (Sec. F). On top of this FuSIC-KBQA design, FUn-FuSIC makes two modifications. First, the logical form generation instruction includes the possibility of unanswerability, and the few shots include examples of the two types of unanswerable questions. The detailed prompt is in the Appendix (Sec. G.1). Secondly, the generated logical form is refined using feedback as before, but this is now adapted for unaswerability. In the rest of this section, we describe feedback for unaswerability (FUn) in detail.
We observe that the first prong of our solution – prompting for unanswerability (PUn) – largely biases the LLM towards one type of error. For $l=$ NK questions, LLM with PUn mostly generates incorrect logical forms instead of generating NK. The essence of our second prong, FUn, is to assess the LLM’s confidence about any logical form being the correct one for the question. If this confidence is low, then FUn outputs $l=$ NK and $a=$ NA. If there is sufficient confidence in a specific logical form $\hat{l}$, then FUn outputs $l=\hat{l}$. If $\hat{l}$ upon execution returns a non-empty answer $\hat{a}$, then FUn outputs $a=\hat{a}$, else the output is $a=$ NA.
We illustrate flow of FUn using examples in Fig. 1. The complete algorithm is in the Appendix (Sec. 2). FUn assesses confidence in a specific logical form using a suite of checks – syntactic, semantic, and execution-based. Some of these are performed using rules defined for the KB, while others involving uncertainty are performed using LLM prompting. FUn goes over $n$ iterations. In the $i^{th}$ iteration, it generates a new logical form $l^{(i)}$ based on feedback received from checks in earlier iterations. $l^{(i)}$ goes through a sequence of checks. If it passes all checks, FUn outputs $l=l^{(i)}$. If $l^{(i)}$ fails a certain check, it is rejected. On the other hand, if it fails some combination of uncertain checks, it is added to the candidate logical form set $L$. The error message from each failed check is appended to the prompt for the next iteration.
If no clear winner emerges at the end of $n$ iterations, FUn assesses the LLM’s confidence about $L$ as a whole, and also about each logical form in $L$. For this, we propose self consistency with unanswerability (scUn). Let $L_{ne}\subset L$ indicate candidate logical forms that produce non-empty answers upon KB execution. Let $a^{*}$ denote the most popular answer (ties being broken at random) and $L_{ne}(a^{*})\subset L_{ne}$ denote logical forms that produce $a^{*}$. If $|L_{ne}(a^{*})|>t$, where $t$ is a threshold, scUn decides that $a=a^{*}$. LLM prompting then selects the most suitable candidate $l^{*}$ from $L_{ne}(a^{*})$, and FUn outputs ($l=l^{*}$, $a=a^{*}$).
If no consensus emerges from $L_{ne}$, scUn turns to its complement set, $L_{e}=L-L_{ne}$, having empty answers. These represent candidates that agree on $a=$ NA. If $L_{e}$ is not empty, again LLM prompting identifies the most suitable candidate $l^{*}$ from $L_{e}$, and FUn outputs ($l=l^{*}$, $a=$ NA). If on the other hand, $L_{ne}$ is empty, FUn outputs ($l=$ NK, $a=$ NA).
Next, we discuss the different categories of checks. The high level goal is to prune candidate logical forms without knowledge of the gold logical form, the gold answer or even whether the question is unanswerable or not. The specific prompts for the checks and the detailed structures of the feedback containing the error messages are described in the Appendix (Sec. G.2).
(C1) Syntax Error:
The KB engine returns a syntax error upon executing the logical form. This is also a part of EGF in FuSIC-KBQA. Note that this is a certain check – meaning that any logical form that fails this check cannot, with certainty, be the right logical form.
(C2) KB Inconsistency:
The logical form is inconsistent with the KB schema. This can happen in different ways. (C2a) Incompatibility in types: The logical form contains a variable and a connecting relation whose types are incompatible. This error is also checked in FuSIC-KBQA EGF. (C2b) Schema hallucinations: The logical form contains schema elements (types, relations, entities) not in the target KB. (C2c) Type-casting errors: Literals are not correctly type-cast for the target KB, e.g. numeric literals as float for Freebase. All of these are certain checks, and are implemented using rules defined using the logical form and the KB.
(C3) Question Logical form Disagreement:
The generated logical form is syntactically correct and consistent with the target KB but is not semantically equivalent to the original natural language question. This is not a certain check for two reasons. First, different types of ambiguities may exist in the question, ruling out the existence of a semantically equivalent logical form in a specific KB. For example, the question may ask for a person from a country, when the KB only contains the relations born in and works in between these types. Secondly, this check requires a learning based classifier, inevitably having some probability of error.
FUn performs this equivalence check using a novel multi-stage LLM pipeline. (i) The variable names in the logical form are first naturalized, preserving semantics and considering the natural language question, e.g. by replacing ’?x’ with ’?actor’. (ii) The naturalized logical form is back-translated into a natural language question. (iii) The back-translated natural language questions is checked for semantic equivalence with the original natural language question. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. More details about few-shot construction are in the Appendix (Sec. G.3).
(C4) Answer Incompatibility:
If the logical form is syntactically correct and consistent with the target KB, it returns some answer upon execution. This answer may not be compatible with the original question. (C4a) The entity returned as the answer is already contained in the logical form (and therefore mentioned in the question). (C4b) The answer contains at least one entity with an intermediate (compound value) type, and therefore does not correspond to any real-world entity. (C4c) The returned answer is empty. All of these checks can be implemented using rules defined over the logical form, the answer and the KB. Note that while the first two are certain checks, the last is not. In the presence of unaswerability, a question may have $a=$ NA as the gold standard.
We now discuss how FUn identifies candidate logical forms based on these checks and assesses the LLM’s confidence about these.
Identifying Candidate Logical Forms:
In the absence of a clear winner (i.e., a logical form that passes all checks), FUn constructs a candidate set $L$ containing logical forms that are flawed, but not fatally so. Logical forms that fail C1, C2, C4a, or C4b are discarded. Those that fail C3 but pass C4c (have non empty answer but disagree with the question), or fail C4c but pass C3 (have empty answer but are equivalent to the original question) are included in $L$.
Self Consistency for Unanswerability:
Consistency breaks down when all the reasoning paths have low probability, leading to a diverse set of answers with little consensus. For our candidate set $L$ with potentially flawed logical forms, the most popular answer $a^{*}$ may have very few votes. Self consistency for unanswerability (scUn) returns the most popular answer $a^{*}$, only if it receives at least $t$ votes out of $n$. We use $t=\frac{n}{2}$. Otherwise, it returns no consistency.
When there is consistency, scUn outputs the most consistent answer $a^{*}$, empty or non-empty. However, we additionally need to identify the most likely logical form with answer $a^{*}$. When $L(a^{*})\subset L$ has more than one logical form, we use LLM prompting to select the most likely $l^{*}\in L(a^{*})$. Details are in the Appendix.(Sec. A.2)
Adapting FUn for Answerable KBQA:
Though motivated by unanswerability, the core idea of iterative generation of candidate logical forms with feedback from error checks, followed by consistency check is applicable more broadly. To adapt this for answerable KBQA, we make the following simple modifications. (a) Logical forms that fail C4c (empty answer) are not included in the candidate set $L$. (b) Instead of scUn, we fall back on regular self consistency for $L$, so that a consistent answer is always returned. These imply that ($l=l^{*}$, $a=a^{*}$) is the only possible outcome.
4 Experiments
Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Overall Answerable Unanswerable Overall Answerable Unanswerable
F1(L) EM-s F1 EM-s F1(L) F1(R) EM-s F1(L) EM-s F1 EM-s F1(L) F1(R) EM-s
RetinaQA 58.4 42.2 28.7 26.0 88.0 84.8 58.4 49.7 35.8 18.7 15.2 80.7 78.7 56.4
FuSIC-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
FUn-FuSIC 76.1 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability.
We now present the experimental evaluation of FUn-FuSIC. First, we aim to address the following research questions for the few-shot transfer setting for KBQA with unanswerability. (R1) How does FUn-FuSIC compare against SoTA models for KBQA with unanswerability, adapted for the few-shot setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) To what extent do the different components of FUn-FuSIC contribute to its performance? In addition, we ask the following research question for the answerable KBQA setting. (R4) How does FUn-FuSIC compare against SoTA few-shot transfer (answerable) KBQA models?
4.1 Experimental Setup
Datasets:
[…]
We create our target test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from the GrailQAbility and GraphQAbility test sets. Additionally, we create the target few shots by selecting 50 answerable and 50 unanswerable questions uniformly at random from the GrailQAbility dev set and the GraphQAbility train set respectively. Of the answerable questions, schema level unanswerable forms 67% in GrailQAbility and 75% in GraphQAbility. The source WebQSP training set has 2858 questions, all of which are answerable.
Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Schema Level Data Level Schema Level Data Level
F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
RetinaQA 94.1 90.9 79.4 76.3 72.9 14.1 83.2 82.0 72.3 73.7 72.7 12.1
FuSIC-U 85.4 80.6 30.9 86.0 80.0 49.4 86.6 82.6 19.0 83.3 83.3 51.5
FUn-FuSIC 85.8 81.2 70.9 83.8 77.6 36.5 92.4 87.5 75.6 80.3 80.3 34.8
Table 2: Performance of various models for different categories of unanswerable questions in two datasets.
Models for comparison:
There are no existing few-shot transfer KBQA models for unanswerability. RetinaQA Faldu et al. (2024) is the SoTA in-domain model for KBQA with unanswerability. It has been shown to significantly outperform various unanswerability adaptations of all other SoTA in-domain KBQA models, such as RnG-KBQA Ye et al. (2022), TIARA Shu et al. (2022), and Pangu Gu et al. (2023), across different categories of unanswerability. We use the available code for RetinaQA. [^2] https://github.com/dair-iitd/RetinaQA We train all components on WebQSP, using as dev set the corresponding target domain dev set. In the absence of unanswerable questions for training, RetinaQA uses a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this. More details are in the Appendix (Sec. I).
FuSIC-KBQA Patidar et al. (2024) is the SoTA model for few-transfer for KBQA with only answerable questions. It comprehensively outperforms a host of supervised and LLM-equipped KBQA models, such as TIARA Shu et al. (2022), Pangu Gu et al. (2023), KB-Binder Li et al. (2023), gf-LLM Shu and Yu (2024), adapted for few-shot transfer. The logical form generation prompt of FuSIC-KBQA is adapted for unanswerability in the instruction and also the few-shots, which now contain unanswerable exemplars. Specifically, we use 5 shots, of which 2 are answerable, 1 is data-related unanswerable, 1 is schema-related unanswerable, and 1 has missing mentioned entity. This adaption also eliminates execution-guided feedback (EGF) from FuSIC-KBQA, which, being based entirely on empty answer check, is unusable for unanswerability. We name this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, where FUn is removed.
Recall that FUn-FuSIC has two mechanisms for addressing unanswerability via LLM prompting. The first is providing few-shots with unanswerability for logical form generation (PUn). The second is using $n$ iterations of FUn. For a fair comparison, all LLM equipped models should have the same LLM “budget” – equivalently, cumulative prompt size. For this reason, we use FUn-FuSIC with zero-shot generation and 4 iterations of FUn, which roughly equals the prompt length of 5-shot FUn-FuSIC-U.
We use gpt-4-0613 as the LLM for both FUn-FuSIC and FuSIC-KBQA-U. The FUn-FuSIC framework allows flexible use of supervised retrievers. For meaningful comparison, we adapt and use RetinaQA as the retriever for both FUn-FuSIC and FuSIC-KBQA-U. More details about the retriever along with details of compute infrastructure are in the Appendix (Sec. H).
Evaluation Measures:
Following existing literature Patidar et al. (2023); Faldu et al. (2024), we evaluate answers using regular and lenient F1 (F1(R) and F1(L)). F1(L) additionally gives credit for the original answer for the complete KB.
For logical form evaluation, the existing evaluation measure (EM) Ye et al. (2022) works only for s-expression as the language. Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL and RetinaQA in s-expression, we propose a new measure (EM-s) that works for both.
EM-s considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answer sets upon KB execution. Since this is a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable, and found $>98\%$ agreement.
More details are in the Appendix (Sec. C).
Answerable Unanswerable
F1 EM-s F1(L) F1(R) EM-s
FUn-FuSIC 57.0 46.0 92.0 90.0 62.0
sc 63.0 52.0 66.0 64.0 36.0
w/o syntax 57.0 46.0 92.0 90.0 62.0
w/o kb-inc 51.7 42.0 92.0 88.0 44.0
w/o q-lf 47.7 38.0 55.8 48.0 10.0
w/o ans-inc 55.0 44.0 92.0 90.0 62.0
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on WebQSP $\rightarrow$ GraphQAbility. sc denotes replacing self consistency for unanswerability with self consistency. Other rows remove checks for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
4.2 Unanswerability Setting
We first address research question R1. Performances of the different models for the few-shot transfer setting with unanswerability are recorded in Tab. 1. We make the following observations. (a) For answerable questions, FuSIC-U and FUn-FuSIC significantly outperform RetinaQA. Across the two datasets, FUn-FuSIC mostly performs better than FuSIC-U (both F1 and EM-s). (b) For unanswerable questions, FuSIC-U performs well in terms of the F1 scores, but not EM-s, indicating empty answers with incorrect logical forms. RetinaQA performs well in terms of F1 scores and EM-s. FUn-FuSIC performs slightly worse than RetinaQA for WebQSP $\rightarrow$ GrailQAbility, but significantly better for WebQSP $\rightarrow$ GraphQAbility. (c) Overall, FUn-FuSIC significantly outperforms both baselines in terms of EM-s, and is slightly below FuSIC-U in terms of F1(L) for WebQSP $\rightarrow$ GrailQAbility. This establishes usefulness of FUn for few-shot transfer KBQA with unanswerability.
We next briefly address research question R2.
Performance of different models for different categories of unanswerability are recorded in Tab. 2. While all models struggle with data-level unanswerability (EM-s), RetinaQA performs the worst and FuSIC-U is the best. For schema-level unanswerability, FUn-FuSIC is the best in one dataset, while RetinaQA is the best in the other. This indicates that more work is needed for robustness across unanswerability categories, and simultaneous use of PUn (prompting) and FUn (feedback) is the most promising approach.
We next address research question R3. Tab. 3 records performance for different ablations of FUn-FuSIC. The biggest benefit, for both answerable and unanswerable questions, comes from question logical form disagreement feedback. Feedback on KB inconsistency and answer incompatibility also make significant contributions to the performance.
Model WebQSP $\rightarrow$ WebQSP $\rightarrow$
GrailQA-Tech GraphQA-Pop
FuSIC 70.8 52.3
FUn-FuSIC(sc) 73.6 67.0
FuSIC-U 62.6 43.4
FUn-FuSIC(scUn) 71.2 65.0
Table 4: Performance using F1 of different models for few-shot KBQA transfer with only answerable questions. The models in the first two rows assume knowledge of answerability of all questions, while the those in the next two do not.
4.3 Answerable Setting
We now address answerable-only KBQA and research question R4. We use two of the datasets used in existing literature for this setting, including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop) Patidar et al. (2024) [^3] https://github.com/dair-iitd/FuSIC-KBQA/. For consistency with existing literature, here we use TIARA Shu et al. (2022) as the retriever for all models.
This setting can be evaluated in two ways. In the first setting, the models are aware that all questions are answerable. This is the setting studied in the KBQA literature, and FuSIC-KBQA is the established SoTA model outperforming a host of supervised and LLM-based models adapted for the task Patidar et al. (2024). We appropriately adapt FUn-FuSIC as discussed earlier in Sec. 3.2.
The first two rows in Tab. 4 record performance for this setting. We see that FUn-FuSIC with self-consistency significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This establishes the usefulness of the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks.
In the more practical second setting, which has not been studied before, although all questions are answerable, the models are not aware of this and make predictions assuming unanswerability. For this we evaluate FuSIC-U and FUn-FuSIC as in Sec. 4.2. The bottom two rows of Tab. 4 record their performance. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability do not exist.
4.4 Error Analysis
We have performed detailed error analysis for FUn-FuSIC. For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions with EM-s $<1$ (46.2% of questions). There are two high-level error categories. (E1) Retrieval Error: At least one necessary KB element (class, relation, entity) is missing in the retrieved KB context (23.4% of questions). (E2) Generation Error: Despite correct retrieval, the generated logical form is incorrect (22.8% of questions). Generation errors can be further sub-categorized. (E2a) The gold logical form is NK, but some (non-NK) logical form is predicted (8.4% of questions). (E2b) A (non-NK) gold logical form exists, but NK is predicted, (4.6% of questions). (E2c) The gold and the predicted logical forms are non-NK, but different (9.8% of questions). We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms in a complete KB. Incompleteness eliminates one of these, so that the gold logical form is NK, while the model identifies the other still existing non-NK logical form. (2) FUn is unable to generate the gold logical form within its maximum number of iterations ($n=4$). (3) FUn generates the gold logical form but it is rejected incorrectly by the question logical form disagreement checker.
5 Conclusions
For robust, low-resource KBQA, we have proposed the practical task of few-shot transfer learning with unanswerability. We have proposed FUn-FuSIC that adapts the existing SoTA model that assumes answerability by (i) candidate logical form generation by iterative LLM prompting with feedback from diverse error checks – including a novel back-translation based equivalence check, and (ii) assessing confidence in the candidates by adapting self consistency to detect answerability and different categories of unanswerability. Using newly created datasets, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable KBQA. Our error analysis suggests that performing well across all categories of unanswerability in the few-shot transfer setting is still a challenge and should be a focus of further research.
6 Limitations
Since LLM inference involves randomness, ideally experiments should be repeated for multiple runs and results should report averages and error bars. Unfortunately, we were not able to do this due to the prohibitive cost of GPT-4, and our results are based on single runs.
While GPT-4 is currently the best performing LLM, it is proprietary as well as expensive. Ideally, evaluation should involve open-source freely accessible LLMs as well. We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by an open LLM. Earlier research has shown that models using Mistral are still able to outperform fully supervised models for answerable few-shot transfer Patidar et al. (2024). Whether this trend holds for the unanswerable setting is an open question. That said, following current trends, we expect the ability of open LLMs to steadily improve in the coming years.
7 Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge. However, can incorrectly inferring unaswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk? While we cannot imagine such a risk at the present time, this may require more thought. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
Appendix A Appendix
A.1 KBQA Problem Formulations
We begin by defining the few-shot transfer learning setting for Knowledge Base Question Answering (KBQA) with unanswerability, and then present our solution for it.
A.1.1 KBQA:
A Knowledge Base $G$ consists of a schema with data stored under it. The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. We are given a target Knowledge Base $G^{t}$ (consisting of entity types $T^{t}$, relations $R^{t}$, entities $E^{t}$ and facts $F^{t}$) and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns a non-empty answer $A^{t}$ for the question $q^{t}$.
A.1.2 Few Shot Transfer Learning for KBQA:
In few-shot transfer learning, we are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms in the target domain. In addition, we are given access to a related source domain. This has a source knowledge base $G^{s}$ consisting of its own types $T^{s}$, relations $R^{s}$, entities $E^{s}$ and facts $F^{s}$, and a much larger source training set $D^{s}$ containing thousands of labeled training examples of questions and corresponding logical forms. The source and target tasks may differ significantly. First, the data and schema of the knowledge bases $G^{t}$ and $G^{s}$ and the domains they cover may be different. Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
A.1.3 KBQA with unanswerability:
A natural language question q is said to be answerable for a Knowledge Base $G^{t}$ if it has a corresponding logical form $l$ which when executed over $G^{t}$ returns a non-empty answer $A$. In contrast, a question $q$ is unanswerable for $G^{t}$, if it either (a) does not have a corresponding logical form that is valid for $G^{t}$, or (b) it has a valid logical form $l$ for $G^{t}$, but which on executing returns an empty answer. The first case indicates that $G^{t}$ is missing some schema element necessary for capturing the semantics for $q$. The second case indicates that the types $T^{s}$, relations $R^{s}$ is sufficient for $q$, but $G^{t}$ is missing some necessary data elements for answering it. In the KBQA with unanswerability task, given a question $q$, if it is answerable, the model needs to output the corresponding logical form $l$. and the non-empty answer $A$ entailed by it, and if it is unanswerable, the model either needs to output $NK$ (meaning No Knowledge) for the logical form, or a valid logical form $l$ with $NA$ (meaning No Answer) as the answer.
Different Types of Unanswerability:
Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer. Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Entity: The topic entity specified in the question is missing from the KB. Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediated entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB. However, the (subject, relation, object) path is not connected in the KB.
A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability:
This setting is a generalisation of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$. The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form. The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
A.2 Algorithm
Figure 2: Feedback Algorithm for Robustness in KBQA
1:Given: KB-Engine, KB-Ontology, qn $\triangleright$ qn is natural language query
[…]
3:$L\leftarrow[]$
4:for $i=0$ to $N$ do
5:     $lf\leftarrow LLM(prompts)$
6:     $syntax\_error\leftarrow KB-Engine(lf)$
7:     if $syntax\_error\neq\text{null}$ then
8:         $syntax\_prompt\leftarrow SynPrompt(lf,syntax\_error)$
9:         $prompts.append(syntax\_prompt)$
10:         continue
11:     end if
12:     $kb\_consistency\_error\leftarrow CheckTypes(lf,KB-Ontology)$
13:     if $kb\_consistency\_error\neq\text{null}$ then
14:         $kb\_consistency\_prompt\leftarrow LFPrompt(lf,kb\_consistency\_error)$
15:         $prompts.append(kb\_consistency\_prompt)$
16:         continue
17:     end if
18:     $qlf\_error\leftarrow CheckMeaning(lf,qn)$
19:     $ans\_consistency\_error\leftarrow CheckAnswer(KB-Engine(lf),lf.nodes)$
20:     if $qlf\_error\neq\text{null}$ then
21:         $qlf\_prompt\leftarrow NLPrompt(lf,qlf\_error)$
22:         $prompts.append(qlf\_prompt)$
23:         if $ans\_consistency\_error$ = null then
24:              $L.append(lf)$
25:         end if
26:         continue
27:     end if
28:     if $ans\_consistency\_error\neq\text{null}$ then
29:         $ans\_consistency\_prompt\leftarrow EGF(lf,egf\_error)$
30:         $prompts.append(ans\_consistency\_prompt)$
[…]
32:              $L.append(lf)$
33:         end if
34:         continue
35:     end if
36:     return $lf$triangleright all checks passed
37:end for
38$\triangleright$ despite N-tries, no lf passes
39:$L_{ne}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)\neq\text{null}]$
40:$L_{e}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)=\text{null}]$
41:$poss\_answers\leftarrow[KB-Engine(lf)\text{ for }lf\text{ in }L_{ne}]$
42:if $scUn(poss\_answers)\neq\text{null}$ then $\triangleright$ No Answer-aware-self-consistency
43:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==scUn(poss\_answers)])$
44:else if $is\_unanswerable$ then
45:     if $L_{e}\neq[]$ then
46:         return $SelectBestLF(qn,L_{e})$
47:     end if
48:     return NK
49:else
50:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==sc(poss\_answers)])$triangleright self-consistency, with tie breaking
51:end if
Algorithm 1 Self-Consistency Algorithm
1:Given: $possible\_answers$
2:Filter out null values from possible_answers:
3:$possible\_answers\leftarrow[p\text{ for }p\text{ in }possible\_answers\text{ if }p\neq\text{null}]$
4:if $possible\_answers=[]$ then
5:     return null
6:end if
7:$MajVote\leftarrow MaxFreqItem(possible\_answers)$
8:return $MajVote$
Algorithm 2 Self-Consistency with Unanswerability Algorithm
1:Given: $possible\_answers$
2:$MajVote\leftarrow self-consistency(possible\_answers)$
3:if $MajVote=\text{null}$ then
4:     return null
5:end if
6:$count\leftarrow Cnt(MajVote,possible\_answers)$
7:if $count>\frac{\text{len}(possible\_answers)}{2}$ then
8:     return $MajVote$
9:else
10:     return null
11:end if
A.3 Working Example of The Algorithm
Question: Where all were the Olympics held before 2024?
Logical Form 1 ?y{ ?x olympics.venue ?y
    ?y olympics.year ?z
    FILTER(?z = 2024)}
Feedback 1 Type: Logical Form Semantic Error Description: ?y is assigned types [’venue’, ’olympics’] by relations [’olympics.venue’, ’olympics.year’]. These types are mutually incompatible.
Logical Form 2 ?y{ ?x olympics.venue ?y
    ?x olympics.year ?z
    FILTER(?z = 2024)}
Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked returns all places where the Olympics were held before 2024. However, the question we answer returns the places where the Olympics were held in 2024.
Logical Form 3 ?y{ ?x olympics.venue ?y
    ?x olympics.year ?z
    FILTER(?z < 2024)}
Feedback 3 Status: Passes All Checks Details: lf3 satisfies syntax, logical form semantic, natural language semantic and execution checks.
Appendix B Ablations
We perform ablations of our best performing model - FUn-FuSIC by removing different types of feedback from the model. We report performance on 100 questions(50 answerable + 50 unanswerable) sampled from the WebQSP $\rightarrow$ GraphQAbility dataset. Refer to 5
Ablation WebQSP $\rightarrow$ GraphQAbility
Overall Ans Unans
F1(L) SM F1 SM F1(L) F1(R) SM
FUn-FuSIC 74.5 54.0 57.0 46.0 92.0 90.0 62.0
sc 64.5 44 63.0 52.0 66.0 64.0 36.0
w/o syntax 74.0 54.0 57.0 46.0 92.0 90.0 62.0
w/o kb-inc 71.8 43.0 51.7 42.0 92.0 88.0 44.0
w/o qf-lf 51.7 24.0 47.7 38.0 55.8 48.0 10.0
w/o ans-inc 73.5 53.0 55.0 44.0 92.0 90.0 62.0
Table 5: Detailed ablation Performance of FUn-FuSIC on WebQSP $\rightarrow$ GraphQAbility. sc denotes replacing self consistency for unaswerability with standard self consistency, while other rows indicate removing check for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
Appendix C EM-s: EM for sparql
As has been observed in Patidar et al. (2023), answer evaluation by itself is not a robust measure for evaluation of KBQA models when the dataset contains unanswerability. Traditional KBQA models that generate s-expressions can be evaluated using EM, which checks for logical form equivalence between two logical forms, since it is possible to compare equivalence between two s-expressions efficiently. However, FUn-FuSIC generates sparql queries instead. Directly comparing program equivalence between two sparql queries is an undecideable problem  [^4] https://users.dcc.uchile.cl/ cgutierr/papers/expPowSPARQL.pdf. Patidar et al. (2024) suggests a semi-automatic strategy for comparison of sparql queries. We propose a completely automatic metric for sparql equivalence check. Two sparql queries are equivalent by the EM-s check if (a) the relations occuring in the two queries are same. (b) the entities occuring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same. Note that the EM-s check is necessary, but not sufficient for two sparql queries to be equivalent.
Since these are a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable and found $>98\%$ agreement.
Appendix D Performance across different categories of unanswerability
We dive deeper into the different categories of unanswerability, as explained in Patidar et al. (2023). There exist 2 broad categories of unanswerability- schema level unanswerability(absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability(absence of facts or intermediate entities of the logical form path on the KB).
We expect that (a) due to the poor ability of supervised models to generalize in transfer learning settings, RetinaQA will be struggle to generate correct logical forms for data level unanswerable questions (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions. However, since it is biased to return incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions. (c) FUn-FuSIC should be able to maintain the performance of FuSIC-KBQA on data level unanswerable questions to a large extent, while significantly improving the performance on schema level unanswerable questions.
Based on performance on the WebQSP $\rightarrow$ GrailQAbility and WebQSP $\rightarrow$ GraphQAbility datasets, we validate the that the trends are indeed as expected.
Appendix E Cost Analysis
FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as U-FuSIC and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16.
The cost for generation of logical form from a prompt with 5 in-context examples is \$0.16.
Thus, the approximate cost for inference of one question by U-FuSIC is \$0.32.
The cost for generation of logical form from a prompt with 0 in-context examples is \$0.04. The cost of checking whether two natural language questions are equivalent or not, using few-shot exemplars and chain of thought prompting is also \$0.04.
The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48. The average cost over 50 randomly sampled questions from the test set is around \$0.34.
Hence, the two models are comparable in terms of cost.
Appendix F FuSIC-KBQA: High Level Overview
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA. FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (b) LLM ReRanking: an LLM is used to select the top-10 classes, top-10 relations and top-5 paths among the top-100 elements as ranked by the supervised retriever, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
Appendix G Prompts
G.1 PUn prompt
Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively. Please do not include any other relations, entities and entity types. Your final sparql can have three scenarios: 1. When you need to just pick from candidate sparql. 2. When you need to extend one of candidate sparql using the candidate relations and entity types. 3. When you will generate a new sparql only using the candidate entities, relations and entity types.For entity type check please use this relation "type.object.type".Do not use entity names in the query. Use specified mids. If it is impossible to construct a query using the provided candidate relations or types, return "NK"Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
NK exemplar Question: the tv episode segments spam fall under what subject? Candidate entities: spam m.04vbm Candidate paths: SELECT DISTINCT ?xWHERE ?x0 ns:tv.tv_segment_performance.segment ns:m.04vbm .?x0 ns:tv.tv_segment_performance.segment ?x .?x ns:type.object.type ns:tv.tv_episode_segment . | ... Candidate entity types: tv.tv_series_episode| tv.tv_episode_segment | ... Candidate relations: tv.tv_series_episode.segments (type:tv.tv_series_episode R type:tv.tv_episode_segment)| tv.tv_subject.tv_programs (type:tv.tv_subject R type:tv.tv_program)|... sparql:NK
Question Prompt Question: which school newspaper deals with the same subject as the onion? Candidate entities: the onion m.0hpsvmv Candidate paths: SELECT DISTINCT ?xWHERE ns:m.0hpsvmv ns:book.newspaper.circulation_areas ?x0 .?x0 ns:periodicals.newspapers ?x .?x ns:type.object.type ns:book.newspaper . |... Candidate entity types: education.school_newspaper| type:book.newspaper... Candidate relations: education.school_newspaper.school (type:education.school_newspaper R type:education.educational_institution) | book.newspaper_issue.newspaper (type:book.newspaper_issue R type:book.newspaper)|... sparql:
G.2 FUN prompt
syntax error feedback Correct the syntax of the following sparql query. Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ... Virtuoso error: word AND not defined
Target KB Disagreement feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query. The assigned relation types by [’computer.computer_emulator.computer’, ’type.object.type computer.computer_peripheral’] are [’computer.computer’, ’computer.computer_peripheral’]. These types are mutually incompatible... Please generate again a different executable sparql using the same context and constraints. DO NOT APOLOGIZE - just return the best you can try.
Question Logical form disagreement feedback The question that you answer is NOT same as what you’ve been asked for! You have answered the question "Which opera productions has Gino Marinuzzi conducted?" but you were asked to answer "what is the name of the premiere opera production conducted by gino marinuzzi?". Please generate again a different executable sparql using the relations, classes and entities provided earlier. DO NOT APOLOGIZE - just return the best you can try.
The next three prompts fall under the answer incompatibility feedback
empty answer feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
intermediate node feedback The generated sparql returns an intermediate type node when executed on the freebase KG. Maybe the answer node is an adjacent node to what we currently query for. Please generate again a different executable sparql using the same context and constraints.
answer node in question feedback The logical form upon execution returns International System of Units, which is not answering the question. Please reconstruct the query using same context and constraints.
G.3 Question Logical Form Agreement Prompt
The few shots provided for verifying question logical form agreement come from the few shots provided in the target domain $D^{t}$. We obtain positive samples from the dataset $D^{t}$ directly, using the questions and gold logical forms. For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over the dataset $D^{t}$. Then we consider those questions for which the predicted logical form is different from the gold logical form.
Firstly, we perform back-translation to obtain natural language question from the logical form
Naturalization of variable names change the sparql query to have variable names representative of what objects they refer to. transform the variable names in this query. Do NOT change the prefix headers and relation names
[…]
We use few-shot LLM prompting to obtain the explanation for why the question and logical form agree or disagree. These few-shots, for obtaining the explanation are dataset independent, and are manually written.
explanation generation prompt Explain why the two questions are different. Question we answer: who all like to eat apple or mango? Question originally asked: what are the people who enjoy both apple and mango? explanation: The question we answer returns people. The question originally asked also returns people. The question we answer finds those people who like eating apple, those people who like eating apple. The question originally asked also finds those people who like eating apple, those people who like eating apple. The question we answer uses logical operator OR. However, the question originally asked uses the logical operator AND Hence, they are different. [total 3 exemplars] Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation:
Some examples of few shots are provided below-
question logical form agreement check Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed atleast 80? explanation:
G.4 Logical Form Selection Prompt
We store the back-translated natural language questions along with each predicted logical form. Here, we prompt the LLM to select the closest back-translated natural language question to the orginal question.
Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director? 1. pred_nl: Which surfing films has Sarah Finn directed the casting for? 2. pred_nl: Which surfing films has Sarah Finn been the casting director for? of the 2 predicted nl questions, which is closest to the original nl question. Even if none is very close, return the one that is semantically closest? pls explain your answer as well
Appendix H Supervised Models training details
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github 555 https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^6] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^7] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^8] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^9] https://github.com/dair-iitd/RetinaQA .
Appendix I Supervised Models inference details
We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths. We might (i) not provide candidate paths (ii) provide candidate paths in GrailQA format (iii) provide candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (b) determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labelled as NK. We choose the optimal value of the threshold to maximise the overall EM-s score over the dev set.

### v2

@@
-Robust Few-shot Transfer Learning for Knowledge Base Question Answering with Unanswerable Questions
+Iterative Repair with Weak Verifiers for Few-shot Transfer in KBQA with Unanswerability
 Abstract
-Real-world KBQA applications require models that are (1) robust – e.g., can differentiate between answerable and unanswerable questions, and (2) low-resource – do not require large training data.
-Towards this goal, we propose the novel task of few-shot transfer for KBQA with unanswerable questions.
-We present FUn-FuSIC that extends the state-of-the-art (SoTA) few-shot transfer model for answerable-only KBQA to handle unanswerability.
-It iteratively prompts an LLM to generate logical forms for the question by providing feedback using a diverse suite of syntactic, semantic and execution guided checks, and adapts self-consistency to assess confidence of the LLM to decide answerability.
-Experiments over newly constructed datasets show that FUn-FuSIC outperforms suitable adaptations of the SoTA model for KBQA with unanswerability, and the SoTA model for answerable-only few-shot-transfer KBQA.
+Real-world applications of KBQA require models to handle unanswerable questions with a limited volume of in-domain labeled training data.
+We propose the novel task of few-shot transfer for KBQA with unanswerable questions and contribute two new datasets for performance evaluation.
+We present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA, the state-of-the-art few-shot transfer model for answerable-only KBQA.
+We first note that FuSIC-KBQA’s iterative repair makes a strong assumption that all questions are answerable.
+As a remedy, we propose Feedback for Unanswerability (FUn), which uses iterative repair using feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerabilty to better assess the answerability of a question.
+Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA for answerable few-shot transfer as well.
+Iterative Repair with Weak Verifiers
+for Few-shot Transfer in KBQA with Unanswerability
+Riya Sawhney, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
 1 Introduction
-The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes a Knowledge Base (KB) and a natural language question as input and requires as output a logical form (or program) that produces an answer upon execution over the KB.
-KBQA has important real-world applications, which require KBQA systems to be low-resource (requiring only a few task-specific labeled examples) and robust.
-One important aspect of robustness is the ability to identify unanswerable questions, which arise naturally in real applications Patidar et al.
+The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes as input a Knowledge Base (KB) and a natural language question, and outputs a logical form (or program) that produces the answer upon execution over the KB.
+KBQA has important real-world applications, which can require systems to be low resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically the ability to identify questions that cannot be answered based on existing KB.
+Traditional supervised models (e.g., Ye et al.
+(2022); Shu et al.
+(2022); Gu et al.
+(2023)) and even recent LLM few-shot in-context learning (FS-ICL) architectures Li et al.
+(2023); Nie et al.
+(2024) fall short in both aspects.
+Limited recent work has addressed these independently – in-domain methods for KBQA with unanswerability trained with large labeled data Patidar et al.
+(2023); Faldu et al.
+(2024), and FuSIC-KBQA for few-shot transfer assuming questions are answerable Patidar et al.
 (2024).
-Traditional supervised models Saxena et al.
+No existing work simultaneously addresses both desiderata in a single KBQA system.
+In response, we propose the novel task of few-shot transfer learning for KBQA with unanswerability.
+Specifically, the target domain may offer only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
+For KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative error guided repair.
+Specifically, multiple iterations of execution-guided feedback are run until a non-empty answer is obtained – this assumes that all questions are answerable.
+For our task, we first design FuSIC-KBQA-U, which, to address Unanswerability, eschews the now-inappropriate repair step, and modifies the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars.
+Unlike in studies for unanswerability in general QA Slobodkin et al.
+(2023), this vanilla adaptation mostly generates incorrect logical forms for unanswerable questions.
+As a remedy, we design a novel solution: FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA).
+The key idea is to modify iterative repair, which earlier relied on a single strong verifier for the logical form’s incorrectness, to rely on a suite of strong and weak verifiers, where strong verifiers identify certain errors, whereas weak verifiers identify potential errors in the current logical form.
+FUn-FuSIC’s verifiers consider both the logical form and the answer.
+For answers, non-emptiness check is now a weak verifier, given potentially unanswerable questions.
+For logical forms, we use strong verifiers to identify obvious syntactic and semantic errors.
+We also propose a novel verifier: non-equivalence of the original question and the back-translation of the logical form.
+This is weak, due to potential errors in back-translation as well as in equivalence classification.
+Using such iterative weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms.
+Simply selecting from this set the logical form with majority answer via self-consistency Wang et al.
+(2023) breaks down in the face of unanswerability.
+We introduce self-consistency for unaswerability, which additionally checks the likelihood of the majority answer, empty or otherwise, to produce the final output.
+For evaluation in our novel task, we create two KBQA transfer datasets with unanswerable questions in the target.
+Our experiments show that FUn-FuSIC comprehensively outperforms different types of SoTA models suitably adapted for this task, including LLM-based and more traditional models.
+We further find that iterative repair of logical forms using weak verifiers holds promise for KBQA in general.
+Using experiments over benchmark datasets for answerable-only few-shot KBQA transfer, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the known SoTA on the task.
+In summary, our specific contributions are as follows.
+(a) We propose the problem of few-shot transfer for KBQA with unanswerability.
+(b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of weak verifiers.
+(c) We create new datasets for the proposed task, which we make public.
+[^1] https://anonymous.4open.science/r/FUn-FuSIC-1704 (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting.
+(e) We also show that for standard answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
+2 Related Work
+In-domain KBQA using supervised models  Saxena et al.
 (2022); Zhang et al.
@@
 (2022); Gu et al.
-(2023) and even recent LLM-based few-shot in-context learning (FS-ICL) architectures Li et al.
+(2023) and using LLM few-shot approaches Li et al.
 (2023); Nie et al.
-(2024) are found wanting in both aspects.
-Recent work has addressed low-resource KBQA in the form of few-shot transfer learning Patidar et al.
-(2024) – the idea is to exploit existing KBQA datasets (source) and transfer them few-shot to a target setting.
-Their FuSIC-KBQA architecture uses supervised KB retrieval followed by generation and refinement of logical forms using LLM few-shot in-context learning.
-A parallel direction addresses unanswerable questions for KBQA Patidar et al.
-(2023).
-Unanswerability arises due to deficiencies in the schema or the data in the KB.
-Faldu et al.
-(2024) proposes RetinaQA – a supervised architecture (with high training data needs) that is robust to both answerable and unanswerable questions.
-However, no existing architecture simultaneously addresses the low-resource and robustness requirements.
-In response, we propose the task of few-shot transfer learning for KBQA in the unaswerability setting.
-Specifically, the target KBQA task offers only a few labeled examples, while the source KBQA task has thousands of labeled examples containing answerable questions only.
-Building a model for this setting could involve extending FuSIC-KBQA for unanswerability or RetinaQA for few-shot transfer.
-Since RetinaQA’s design requires large volumes of labeled training data in the target and it fares poorly in few-shot setting, we build our model on top of FuSIC-KBQA.
-A natural vanilla extension is to modify the LLM prompt for unanswerability and including different types of unanswerable questions in the few shots, as proposed in the broader unanswerable QA literature Slobodkin et al.
-(2023).
-Unfortunately, we find that this largely fails for unanswerable questions that do not have a valid logical form – the simple modification regularly labels these as answerable and outputs incorrect logical forms.
-A second issue is that FuSIC-KBQA prompts LLM to refine the logical form if its execution leads to an empty answer – but this assumes that all questions are answerable and naturally breaks down in the face of unanswerability.
-To handle this, we develop FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA).
-FUn-FuSIC takes the vanilla approach further by detecting when the LLM is very uncertain about the logical form, which is suggestive of the question being unanswerable.
-For this, it employs logical form refinement by providing LLM repeated feedback about potential errors, including syntactic, semantic and execution-related error checks, and a novel strategy of back-translating the logical form to check equivalence with the original question.
-To assess the model’s lack of certainty about the logical form, FUn-FuSIC adapts self-consistency Wang et al.
-(2023) – in the absence of enough certainty, schema-level unanswerability is inferred.
-On the other hand, when a sufficient number of logical forms agree on the answer (empty or non-empty), the LLM is finally prompted to choose the best one, along with inferring answerability (for non-empty answer) or data-related unanswerability (for empty answer).
-For evaluation, we first create two new datasets for few-shot transfer for KBQA with unanswerability.
-Using these, we show that, in aggregate, FUn-FuSIC comprehensively outperforms RetinaQA, when adapted for few-shot transfer, as well as the vanilla extension to FuSIC-KBQA.
-We further demonstrate that the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks, holds promise for KBQA in general.
-Using experiments over existing few-shot transfer datasets for answerable KBQA, we find that the restriction of FUn-FuSIC for the answerable setting outperforms FuSIC-KBQA, which is the SoTA KBQA model in this setting.
-In summary, our specific contributions are as follows.
-(a) We propose the problem of few-shot transfer for KBQA with unanswerability.
-(b) We present FUn-FuSIC that uses iterative feedback for error checks and self-consistency adapted for unaswerability.
-(c) We create new datasets for the proposed task which we make public  [^1] https://anonymous.4open.science/r/FUn-FuSIC-1704.
-(d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting.
-(e) We also show that in the standard answerable setting, FUn-FuSIC outperforms the corresponding SoTA model.
-2 Related Work
-There is a lot of work on in-domain KBQA using supervised models  Ye et al.
-(2022); Chen et al.
-(2021); Shu et al.
-(2022); Gu et al.
-(2023) as well as LLM few-shot approaches Li et al.
-(2023); Nie et al.
-(2024); Shu and Yu (2024).
-These require significant volumes of labeled data to train or select the most appropriate few shots.
-Also in the in-domain setting, unanswerability has been studied for KBQA Patidar et al.
+(2024); Shu and Yu (2024) is well explored in literature.
+These use high volumes of labeled data, either for training or selecting the most relevant few shot exemplars.
+For in-domain KBQA, unanswerability has recently been studied Patidar et al.
 (2023); Faldu et al.
@@
 Patidar et al.
-(2023) created the GrailQAbility dataset with different categories of unanswerability, and showed that answerable-only models, when adapted superficially for unaswerability, are inadequate in different ways.
+(2023) create the GrailQAbility dataset with different categories of unanswerability, and show the inadequacy of superficial adaptations of answerable-only KBQA models.
 RetinaQA Faldu et al.
-(2024) is the SoTA model for unaswerability.
-It learns to discriminate between valid logical forms and those rendered invalid due to schema-level incompleteness, and constructs logical forms top-down using sketch-filling to handle data-level incompleteness.
+(2024) is the SoTA model for KBQA unaswerability.
 However, this also requires large volumes of training data.
-In the transfer setting for KBQA, one formulation makes use of large volumes of unlabeled training data in the target Cao et al.
+For KBQA transfer Cao et al.
 (2022); Ravishankar et al.
+(2022), low-resource was originally not a focus.
+More recently, few-shot transfer for KBQA has been addressed Patidar et al.
 (2024).
-More recently, FuSIC-KBQA has addressed the few-shot transfer setting with only a small volume of labeled training examples in the target Patidar et al.
-(2024).
-This uses a retrieve-then-generate framework using one or more supervised retrievers and LLM few-shot in-context learning based generators.
-The LLM generation also goes through multiple iterations of execution-guided feedback until a non-empty answer is obtained.
-However, the transfer setting has only considered answerable questions.
-LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al.
-(2023), but without feedback for iterative generation.
+Its FuSIC-KBQA model uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair.
+However, this formulation assumes answerability of all questions, as described in Section 1.
+Simple LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al.
+(2023), but without any notion of feedback or iterative repair.
 Other approaches  Shinn et al.
 (2023); Chen et al.
-(2023b) use execution based refinement for programs but do not involve unanswerability.
-In the prompting literature, self-consistency marginalizes over multiple reasoning paths to arrive at a consensus answer Wang et al.
+(2023b) use execution based refinement for program generation but without any notion of unanswerability.
+The extensive recent work on program self-repair using LLMs Olausson et al.
+(2024) also diverges strongly from our few-shot goal.
+It assumes availability of unit tests for every test instance.
+More importantly, as in the answerable KBQA literature, it assumes that a valid program exists for each specification.
+To the best of our knowledge, ours is the first work to apply iterative weak repair for KBQA, and to use iterative repair for unanswerability.
+3 Background & Problem Definition
+Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question $Q$ when executed over different KBs with $\leq 4$ iterations.
+$Q$ is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness).
+FUn iterations are shown using gray blocks, with Strong Verifiers in red and Weak Verifiers in green.
+(V1: Syntax Error omitted for brevity) scUn is shown using the blue block.
+A Knowledge Base (KB) $G$ consists of a schema and data.
+The schema consists of entity types (or classes) $T$ and binary relations $R$ defined over pairs of types.
+The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$.
+Given a target KB $G^{t}$ and a natural language question $q^{t}$, the goal is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ (which is a set in general).
+Target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms.
+A related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
+Following Patidar et al.
+(2023), a question $q$ is answerable for a KB $G$ if it has a corresponding logical form $l$ which when executed over $G$ returns a non-empty answer $A$.
+A question is unanswerable, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$ (data-level unanswerability).
+Schema-level unanswerability arises due to missing types and relations, while missing entities and facts lead to data-level unanswerability.
+More details are in Appendix  A.1.3.
+In KBQA with unanswerability, given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
+We now define few-shot transfer learning for KBQA with unanswerability.
+A target question $q^{t}$ may be answerable or unanswerable due to missing schema or data in the target KB $G^{t}$.
+Target few shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories.
+However, most KBQA datasets contain only answerable questions, we model the source training data $D^{s}$ as containing only answerable questions.
+Now the unanswerability gap between the source and the target also needs to be bridged.
+More details are in Appendix A.1.
+4 Proposed Approach: FUn-FuSIC
+Our proposed architecture FUn-FuSIC adapts FuSIC-KBQA for unaswerability.
+The high-level algorithm is described in Algo.
+1.
+Preserving the retrieve-then-generate framework, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers.
+Each retriever $R_{i}$ is source-trained and then target fine-tuned.
+The retrieval output $r$ of each $R_{i}$ consists of relevant schema elements (types and relations) for $q^{t}$, and data paths emanating from mentioned entities in $q^{t}$.
+The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few shots $D^{t}$.
+More details of the retrieval stage are in the Appendix (Sec.
+A.6).
+Algorithm 1 FUn-FuSIC$(q,G^{t},D^{t},R,V^{s},V^{w},\mathcal{L})$
+1:$r=\{\}$
+[…]
+3:$l=\mbox{PUn}(\mathcal{L},I,q,r,D^{t})$
+4:$(e,l,A,L)=\mbox{FUn}(\mathcal{L},q,l,n,V^{s},V^{w},G^{t})$
+5:if ($\mbox{{\bf if} }(e)\mbox{ {\bf return}}(l^{*},A^{*})$
+6:else $\mbox{{\bf else} {\bf return}}(\mbox{scUn}(q,L,t,\mathcal{L}))$
+FUn-FuSIC makes three modifications to the FuSIC-KBQA architecture.
+First, the LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions.
+However, this simple approach is error-prone.
+So, we bias the instruction towards one type of error.
+Specifically, when uncertain about answerability of the question, Prompting for Unanswerability (PUn) (line 3) instructs $\mathcal{L}$ to generate a (possibly incorrect) logical form instead of $l=\texttt{NK}$.
+The detailed prompt $I$ is in the Appendix (Sec.
+A.8).
+We now come to the more significant modifications.
+First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unaswerability.
+This iterative repair, which we name Feedback for Unaswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed.
+For this, we introduce self-consistency for Unaswerability (scUn).
+scUn checks the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output.
+In the rest of this section, we describe FUn and then scUn.
+Fig.
+1 illustrates flow of FUn and scUn using examples.
+A real example of FUn execution is in Sec.
+A.13.
+Algorithm 2 $\mbox{FUn}(q,l,n,V^{s},V^{w},G,\mathcal{L})$
+1:$i=0$, $k=0$, $L=\{\}$, $F=\mbox{""}$"
+2:while $++i\leq n$ do
+3:     for $j=1\mbox{ to }k_{1}$ do
+4:         $(e,f)=V^{s}_{j}(l,q,G)$
+5:         $F=\mbox{Append}(F,f)$
+6:         $\mbox{{\bf if} }(!e)\mbox{ {\bf break}}$
+7:     end for
+8:     for $j=1\mbox{ to }k_{2}$ do
+9:         $(e,f)=V^{w}_{j}(l,q,G)$
+10:         $F=\mbox{Append}(F,f)$
+11:         if ($\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l\}$; $k++$
+12:     end for
+13:     $l=Gen(\mathcal{L},I,q,F)$
+14:end while
+15:if ($\mbox{{\bf if} }(k=k_{2})\mbox{ {\bf return}}(\mbox{True},l,\mbox{Exec}(l,G),L)$
+16:else $\mbox{{\bf else} {\bf return}}(\mbox{False},l,\{\},L)$
+The FUn algorithm is described in Algo.
+2 Starting with the initial logical form $l^{(0)}$, FUn performs at most $n$ verify-and-repair iterations to create a candidate set $L$ of probable logical forms.
+The example shows 3 FUn iterations for KB1 and KB2, and 2 iterations for KB3.
+In the $i^{th}$ iteration, FUn generates a new logical form $l^{(i)}$ by prompting $\mathcal{L}$ using $q^{t}$ and feedback $F$ received from checks in all previous iterations (line 13).
+$l^{(i)}$ goes through a sequence of verifications.
+FUn uses two sets of verifiers.
+The strong verifiers $V^{s}$ are guaranteed to be correct, while the weak verifiers $V^{w}$ are potentially erroneous.
+The specific verifiers that we use in this paper are defined later in the section.
+A template-based feedback string $f$ is appended to the generation prompt for $l^{(i+1)}$ based on the specific verifier that $l^{(i)}$ failed.
+If $l^{(i)}$ fails a strong verifier, it is rejected (line 6).
+In the example, this happens for all three KBs in iteration 1.
+If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$.
+This happens in iteration 3 for KB3.
+Otherwise, if $l^{(i)}$ passes at least one weak verifier but not all, it is added to candidate logical form set $L$ (line 11).
+This happens for iterations 2, 3 and 4 for KB1 and KB2.
+Logical form verifiers:
+FUn uses a suite verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$).
+These may be syntactic, semantic, or execution-based, defined using simple rules or complex LLM functions over $l$, $q$ and $G$.
+Note that unlike unit tests in program synthesis, the verifiers do not have knowledge of the gold logical form, the gold answer or answerability of the question.
+We now briefly describe the specific verifiers that we use for this paper.
+Additional details about the verifiers are in the Appendix (Sec.
+A.8.1 and Sec.
+A.2).
+Note however that FUn is a framework that is equally capable of working with a wholly different suite of verifiers.
+(V1) Syntax Error:
+As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check: a valid logical form cannot have syntax error.
+(V2) KB Inconsistency:
+A logical form $l$ may be inconsistent with the schema of $G$.
+We identify semantic errors of different categories, such as type-incompatibility and schema hallucinations, implemented using rules over $l$ and $G$.
+These are also strong verifiers.
+(V3) Question - Logical Form Disagreement:
+This verifier checks if $l$ is semantically equivalent to the original natural language question $q$.
+In our example, LF3 for KB1 disagrees with Q.
+This is a weak verifier.
+First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities.
+For example, $q$ mentions a person from a country, when $G$ has the relations born in and works in between these types.
+Secondly, this is implemented as a probabilistic classifier that can err.
+We define equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$.
+More details are in Sec.A.2.
+(V4) Answer Inconsistency:
+This verifier executes $l$ over $G$ to obtain an answer $A$ and then checks its compatibility with $q$.
+This may fail for different reasons, such as (V4a) $A$ containing an entity mentioned in $q$, (V4b) $A$ being empty, and others.
+Note that V4a is a strong verifier while V4b is weak, since an empty answer is valid for unanswerable questions (as for LF3 for KB2), but invalid for answerable ones.
+Identifying Candidate Logical Forms:
+Unless some logical form passes all checks and is therefore returned (Algo.
+2 line 6), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so.
+For our specific suite of weak verifiers, $l^{(i)}$ is added to $L$ if it passes one of V4b ($A$ is non-empty) as for LF3 for KB1, or V3 ($l^{i}$ is equivalent to $q$), as for LF3 for KB2.
+Self Consistency for Unanswerability:
+If FUn fails to generate a single confident logical form at the end of $n$ iterations, as for KB1 and KB2, we use scUn to decide if the best candidate $l^{*}\in L$ has sufficient confidence.
+If scUn detects sufficient confidence, as for KB2, it outputs ($l=l^{*}$, $A=A^{*}$), where $A^{*}$ is the answer obtained upon executing $l^{*}$ and may be NA.
+If on the other hand, scUn decides insufficient confidence, as for KB1, it outputs ($l=$ NK, $A=$ NA).
+For identifying the consensus choice from $L$, one possibility is self-consistency Wang et al.
 (2023); Chen et al.
-(2023a).
-This works only when significant probability mass accumulates for at least one answer, which is selected as the winner.
-This breaks down under unaswerability, when reasoning paths to all non-empty answers have low probability.
-3 Problem and Proposed Approach
-Figure 1: Illustration of FUn-FuSIC for a specific question when executed over three different KBs.
-The question is answerable for KB3, but unanswerable for KB1 and KB2, due to schema incompleteness and data incompleteness respectively.
-The iterations as shown as i=1,2,3,4, The consistency block is below these.
-Colored boxes show different checks.
-Additions to the set of candidate logical forms are shown in the gray columns.
-We first briefly cover background on KBQA, unaswerability and few-shot KBQA.
-More details are in the Appendix (Sec.
-A.1).
-We then define the few-shot setting with unanswerability and our proposed approach.
-3.1 Few-shot Transfer Learning for KBQA with Unanswerability
-A Knowledge Base $G$ consists of a schema with data stored under it.
-The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types.
-The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$.
-We are given a target Knowledge Base $G^{t}$ and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns an answer set $a^{t}$.
-We are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms.
-A related source domain has a source knowledge base $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
-The source and target KBs and the distributions of questions and logical forms can be very different.
-In general, a KB may have both answerable and unanswerable questions Patidar et al.
-(2023).
-A question $q$ is answerable for a KB $G$ if it has a corresponding logical form which when executed over $G$ returns a non-empty answer $A$.
-In contrast, a question is unanswerable for $G$, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form for $G$, but which on executing on $G$ returns an empty answer (data-level unanswerability).
-Schema-level unaswerability arises due to missing types (classes) and relations.
-On the other hand, data-level unaswerability arises due to missing entities and facts.
-Absence of any mentioned question entity, however, invalidates the logical form.
-More details are in the Appendix (Sec.
-A.1.3).
-In KBQA with unanswerability, given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
-We now propose the problem of few-shot transfer learning for KBQA with unanswerability.
-Here, the target KB $G^{t}$ is incomplete, and target test questions $q^{t}$ may be answerable or unanswerable due to missing schema or data.
-The few shot examples $D^{t}$ contain both answerable and unanswerable questions of different types.
-Since almost all large KBQA datasets contain only answerable questions, we assume that the source training data $D^{s}$ contains only answerable questions.
-So, in addition to the differences between source and target KBQA tasks in the existing transfer learning setting, the unanswerability gap also needs to be bridged.
-3.2 Proposed Approach: FUn-FuSIC
-Our proposed architecture FUn-FuSIC builds upon and adapts FuSIC-KBQA for unaswerability.
-We preserve its retrieve-then-generate framework where the first stage is KB retrieval with one or more supervised retrievers, which are source-trained and then target fine-tuned.
-The retrieval is LLM reranked to increase target relevance.
-This forms the input to the generative stage, which prompts an LLM to generate the logical form using the retrieval, either in few-shot or zero-shot fashion.
-More details of these stages are in the Appendix (Sec.
-F).
-On top of this FuSIC-KBQA design, FUn-FuSIC makes two modifications.
-First, the logical form generation instruction includes the possibility of unanswerability, and the few shots include examples of the two types of unanswerable questions.
-The detailed prompt is in the Appendix (Sec.
-G.1).
-Secondly, the generated logical form is refined using feedback as before, but this is now adapted for unaswerability.
-In the rest of this section, we describe feedback for unaswerability (FUn) in detail.
-We observe that the first prong of our solution – prompting for unanswerability (PUn) – largely biases the LLM towards one type of error.
-For $l=$ NK questions, LLM with PUn mostly generates incorrect logical forms instead of generating NK.
-The essence of our second prong, FUn, is to assess the LLM’s confidence about any logical form being the correct one for the question.
-If this confidence is low, then FUn outputs $l=$ NK and $a=$ NA.
-If there is sufficient confidence in a specific logical form $\hat{l}$, then FUn outputs $l=\hat{l}$.
-If $\hat{l}$ upon execution returns a non-empty answer $\hat{a}$, then FUn outputs $a=\hat{a}$, else the output is $a=$ NA.
-We illustrate flow of FUn using examples in Fig.
-1.
-The complete algorithm is in the Appendix (Sec.
-2).
-FUn assesses confidence in a specific logical form using a suite of checks – syntactic, semantic, and execution-based.
-Some of these are performed using rules defined for the KB, while others involving uncertainty are performed using LLM prompting.
-FUn goes over $n$ iterations.
-In the $i^{th}$ iteration, it generates a new logical form $l^{(i)}$ based on feedback received from checks in earlier iterations.
-$l^{(i)}$ goes through a sequence of checks.
-If it passes all checks, FUn outputs $l=l^{(i)}$.
-If $l^{(i)}$ fails a certain check, it is rejected.
-On the other hand, if it fails some combination of uncertain checks, it is added to the candidate logical form set $L$.
-The error message from each failed check is appended to the prompt for the next iteration.
-If no clear winner emerges at the end of $n$ iterations, FUn assesses the LLM’s confidence about $L$ as a whole, and also about each logical form in $L$.
-For this, we propose self consistency with unanswerability (scUn).
-Let $L_{ne}\subset L$ indicate candidate logical forms that produce non-empty answers upon KB execution.
-Let $a^{*}$ denote the most popular answer (ties being broken at random) and $L_{ne}(a^{*})\subset L_{ne}$ denote logical forms that produce $a^{*}$.
-If $|L_{ne}(a^{*})|>t$, where $t$ is a threshold, scUn decides that $a=a^{*}$.
-LLM prompting then selects the most suitable candidate $l^{*}$ from $L_{ne}(a^{*})$, and FUn outputs ($l=l^{*}$, $a=a^{*}$).
-If no consensus emerges from $L_{ne}$, scUn turns to its complement set, $L_{e}=L-L_{ne}$, having empty answers.
-These represent candidates that agree on $a=$ NA.
-If $L_{e}$ is not empty, again LLM prompting identifies the most suitable candidate $l^{*}$ from $L_{e}$, and FUn outputs ($l=l^{*}$, $a=$ NA).
-If on the other hand, $L_{ne}$ is empty, FUn outputs ($l=$ NK, $a=$ NA).
-Next, we discuss the different categories of checks.
-The high level goal is to prune candidate logical forms without knowledge of the gold logical form, the gold answer or even whether the question is unanswerable or not.
-The specific prompts for the checks and the detailed structures of the feedback containing the error messages are described in the Appendix (Sec.
-G.2).
-(C1) Syntax Error:
-The KB engine returns a syntax error upon executing the logical form.
-This is also a part of EGF in FuSIC-KBQA.
-Note that this is a certain check – meaning that any logical form that fails this check cannot, with certainty, be the right logical form.
-(C2) KB Inconsistency:
-The logical form is inconsistent with the KB schema.
-This can happen in different ways.
-(C2a) Incompatibility in types: The logical form contains a variable and a connecting relation whose types are incompatible.
-This error is also checked in FuSIC-KBQA EGF.
-(C2b) Schema hallucinations: The logical form contains schema elements (types, relations, entities) not in the target KB.
-(C2c) Type-casting errors: Literals are not correctly type-cast for the target KB, e.g.
-numeric literals as float for Freebase.
-All of these are certain checks, and are implemented using rules defined using the logical form and the KB.
-(C3) Question Logical form Disagreement:
-The generated logical form is syntactically correct and consistent with the target KB but is not semantically equivalent to the original natural language question.
-This is not a certain check for two reasons.
-First, different types of ambiguities may exist in the question, ruling out the existence of a semantically equivalent logical form in a specific KB.
-For example, the question may ask for a person from a country, when the KB only contains the relations born in and works in between these types.
-Secondly, this check requires a learning based classifier, inevitably having some probability of error.
-FUn performs this equivalence check using a novel multi-stage LLM pipeline.
-(i) The variable names in the logical form are first naturalized, preserving semantics and considering the natural language question, e.g.
-by replacing ’?x’ with ’?actor’.
-(ii) The naturalized logical form is back-translated into a natural language question.
-(iii) The back-translated natural language questions is checked for semantic equivalence with the original natural language question.
-The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$.
-More details about few-shot construction are in the Appendix (Sec.
-G.3).
-(C4) Answer Incompatibility:
-If the logical form is syntactically correct and consistent with the target KB, it returns some answer upon execution.
-This answer may not be compatible with the original question.
-(C4a) The entity returned as the answer is already contained in the logical form (and therefore mentioned in the question).
-(C4b) The answer contains at least one entity with an intermediate (compound value) type, and therefore does not correspond to any real-world entity.
-(C4c) The returned answer is empty.
-All of these checks can be implemented using rules defined over the logical form, the answer and the KB.
-Note that while the first two are certain checks, the last is not.
-In the presence of unaswerability, a question may have $a=$ NA as the gold standard.
-We now discuss how FUn identifies candidate logical forms based on these checks and assesses the LLM’s confidence about these.
-Identifying Candidate Logical Forms:
-In the absence of a clear winner (i.e., a logical form that passes all checks), FUn constructs a candidate set $L$ containing logical forms that are flawed, but not fatally so.
-Logical forms that fail C1, C2, C4a, or C4b are discarded.
-Those that fail C3 but pass C4c (have non empty answer but disagree with the question), or fail C4c but pass C3 (have empty answer but are equivalent to the original question) are included in $L$.
-Self Consistency for Unanswerability:
-Consistency breaks down when all the reasoning paths have low probability, leading to a diverse set of answers with little consensus.
-For our candidate set $L$ with potentially flawed logical forms, the most popular answer $a^{*}$ may have very few votes.
-Self consistency for unanswerability (scUn) returns the most popular answer $a^{*}$, only if it receives at least $t$ votes out of $n$.
-We use $t=\frac{n}{2}$.
-Otherwise, it returns no consistency.
-When there is consistency, scUn outputs the most consistent answer $a^{*}$, empty or non-empty.
-However, we additionally need to identify the most likely logical form with answer $a^{*}$.
-When $L(a^{*})\subset L$ has more than one logical form, we use LLM prompting to select the most likely $l^{*}\in L(a^{*})$.
-Details are in the Appendix.(Sec.
-A.2)
-Adapting FUn for Answerable KBQA:
-Though motivated by unanswerability, the core idea of iterative generation of candidate logical forms with feedback from error checks, followed by consistency check is applicable more broadly.
-To adapt this for answerable KBQA, we make the following simple modifications.
-(a) Logical forms that fail C4c (empty answer) are not included in the candidate set $L$.
-(b) Instead of scUn, we fall back on regular self consistency for $L$, so that a consistent answer is always returned.
-These imply that ($l=l^{*}$, $a=a^{*}$) is the only possible outcome.
+(2023a) i.e., obtaining the answer for each $l\in L$, and returning those with the most common answer.
+Self-consistency requires some answer to accumulate enough probability by aggregation over reasoning paths.
+However, for unanswerable questions, no single answer accumulates sufficient probability, and self-consistency returns some low probability answer.
+To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$).
+If so, scUn uses LLM prompting to select the most appropriate supporting logical form $l^{*}\in L$ considering $q$, and outputs ($l=l^{*}$, $A=A^{*}$).
+However, for KB1 in the example, the 3 logical forms among the candidates have 3 different answers, and therefore ($\lfloor\frac{|L|}{2}\rfloor=1$) no consensus emerges for non-empty answers.
+Then scUn considers logical forms from $L$ that agree on $A=$ NA.
+If there are multiple such candidates, scUn selects the most suitable candidate $l^{*}$, again using LLM prompting, and outputs ($l=l^{*}$, $A=$ NA).
+If there is no such candidate, scUn outputs ($l=$ NK, $A=$ NA).
+For KB2 in the example, scUn selects LF3 – the only logical form with empty answer.
+Further details on scUn are in the Appendix (Sec.
+A.14).
+.
 5 Experiments
@@
 Overall Answerable Unanswerable Overall Answerable Unanswerable
-F1(L) EM-s F1 EM-s F1(L) F1(R) EM-s F1(L) EM-s F1 EM-s F1(L) F1(R) EM-s
+F1 EM-s F1 EM-s F1(L) F1(R) EM-s F1 EM-s F1 EM-s F1(L) F1(R) EM-s
 RetinaQA 58.4 42.2 28.7 26.0 88.0 84.8 58.4 49.7 35.8 18.7 15.2 80.7 78.7 56.4
-FuSIC-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
-FUn-FuSIC 76.1 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
+Pangu 54.5 43.8 31.23 29.6 83.8 80.4 58.0 53.4 33.0 30.3 26.4 76.5 74.8 39.6
+FuSIC-KBQA-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
+KB-Binder 43.7 33.0 19.5 16.5 67.9 66.5 49.5 44.3 36.1 27.5 21.6 61.0 61.0 50.7
+FUn-FuSIC 76.6 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
 Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability.
-We now present the experimental evaluation of FUn-FuSIC.
-First, we aim to address the following research questions for the few-shot transfer setting for KBQA with unanswerability.
-(R1) How does FUn-FuSIC compare against SoTA models for KBQA with unanswerability, adapted for the few-shot setting?
+We now present experimental evaluation of FUn-FuSIC.
+For few-shot KBQA transfer with unanswerability, we address the following research questions.
+(R1) How does FUn-FuSIC compare against SoTA models for KBQA, suitably adapted for this setting?
 (R2) How does FUn-FuSIC perform across different categories of unanswerability?
-(R3) To what extent do the different components of FUn-FuSIC contribute to its performance?
-In addition, we ask the following research question for the answerable KBQA setting.
-(R4) How does FUn-FuSIC compare against SoTA few-shot transfer (answerable) KBQA models?
+(R3) How do the different components of FUn-FuSIC contribute to its performance?
+In addition, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
 5.1 Experimental Setup
 Datasets:
+For few-shot KBQA transfer, available datasets have only answerable questions Patidar et al.
+(2024), while our task needs the target dataset to contain unanswerable questions as well.
+So we construct our own transfer datasets, augmenting existing ones.
+GrailQAbility is the only available KBQA dataset with unanswerable questions Patidar et al.
+(2023).
+This was carefully designed starting from GrailQA Gu et al.
+(2021), which has only answerable questions, by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries.
+We use this as one of our targets.
+GraphQA Su et al.
+(2016) is another popular KBQA dataset.
+This has the same back-end KB (Freebase) as GrailQA.
+We create our second target dataset using GraphQA, by replacing its KB with the modified KB in GrailQAbility.
+This introduces unanswerability into GraphQA questions and we label these appropriately as schema-level or data-level unanswerable.
+We name this dataset GraphQAbility.
+We create the two test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from their test sets.
+We create few shots by selecting 100 questions (50 answerable and 50 unanswerable) uniformly at random from the dev set and train set respectively.
 […]
-We create our target test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from the GrailQAbility and GraphQAbility test sets.
-Additionally, we create the target few shots by selecting 50 answerable and 50 unanswerable questions uniformly at random from the GrailQAbility dev set and the GraphQAbility train set respectively.
-Of the answerable questions, schema level unanswerable forms 67% in GrailQAbility and 75% in GraphQAbility.
-The source WebQSP training set has 2858 questions, all of which are answerable.
+WebQSP training set has 2,858 real user questions, which are manually annotated with logical forms.
+This is quite different from GraphQA and GrailQA which contain algorithmically generated logical forms, verbalized by crowd-workers.
+Because of this difference in nature of questions, these are challenging transfer datasets, beyond the unaswerability gap.
 Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
@@
 RetinaQA 94.1 90.9 79.4 76.3 72.9 14.1 83.2 82.0 72.3 73.7 72.7 12.1
+Pangu 91.1 87.9 87.9 69.6 65.9 00.0 77.3 74.4 74.4 73.3 72.7 00.0
 FuSIC-U 85.4 80.6 30.9 86.0 80.0 49.4 86.6 82.6 19.0 83.3 83.3 51.5
+KB-Binder 75.1 73.9 70.1 53.1 51.5 09.5 67.0 65.9 60.9 41.2 41.2 06.8
 FUn-FuSIC 85.8 81.2 70.9 83.8 77.6 36.5 92.4 87.5 75.6 80.3 80.3 34.8
-Table 2: Performance of various models for different categories of unanswerable questions in two datasets.
+Table 2: Model performance for categories of unanswerable questions. FuSIC-U is short hand for FuSIC-KBQA-U.
 Models for comparison:
-There are no existing few-shot transfer KBQA models for unanswerability.
-RetinaQA Faldu et al.
-(2024) is the SoTA in-domain model for KBQA with unanswerability.
-It has been shown to significantly outperform various unanswerability adaptations of all other SoTA in-domain KBQA models, such as RnG-KBQA Ye et al.
-(2022), TIARA Shu et al.
-(2022), and Pangu Gu et al.
-(2023), across different categories of unanswerability.
-We use the available code for RetinaQA.
-[^2] https://github.com/dair-iitd/RetinaQA We train all components on WebQSP, using as dev set the corresponding target domain dev set.
-In the absence of unanswerable questions for training, RetinaQA uses a threshold fine-tuned on a dev set to detect schema-level unanswerability.
-We again use the target dev sets for this.
-More details are in the Appendix (Sec.
-I).
-FuSIC-KBQA Patidar et al.
-(2024) is the SoTA model for few-transfer for KBQA with only answerable questions.
-It comprehensively outperforms a host of supervised and LLM-equipped KBQA models, such as TIARA Shu et al.
-(2022), Pangu Gu et al.
-(2023), KB-Binder Li et al.
-(2023), gf-LLM Shu and Yu (2024), adapted for few-shot transfer.
-The logical form generation prompt of FuSIC-KBQA is adapted for unanswerability in the instruction and also the few-shots, which now contain unanswerable exemplars.
-Specifically, we use 5 shots, of which 2 are answerable, 1 is data-related unanswerable, 1 is schema-related unanswerable, and 1 has missing mentioned entity.
-This adaption also eliminates execution-guided feedback (EGF) from FuSIC-KBQA, which, being based entirely on empty answer check, is unusable for unanswerability.
-We name this model FuSIC-KBQA-U.
-Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, where FUn is removed.
-Recall that FUn-FuSIC has two mechanisms for addressing unanswerability via LLM prompting.
-The first is providing few-shots with unanswerability for logical form generation (PUn).
-The second is using $n$ iterations of FUn.
-For a fair comparison, all LLM equipped models should have the same LLM “budget” – equivalently, cumulative prompt size.
-For this reason, we use FUn-FuSIC with zero-shot generation and 4 iterations of FUn, which roughly equals the prompt length of 5-shot FUn-FuSIC-U.
-We use gpt-4-0613 as the LLM for both FUn-FuSIC and FuSIC-KBQA-U.
-The FUn-FuSIC framework allows flexible use of supervised retrievers.
-For meaningful comparison, we adapt and use RetinaQA as the retriever for both FUn-FuSIC and FuSIC-KBQA-U.
-More details about the retriever along with details of compute infrastructure are in the Appendix (Sec.
-H).
+As few-shot transfer for KBQA with unanswerability is a novel task, there are no existing baselines.
+For in-domain KBQA with unanswerability, RetinaQA Faldu et al.
+(2024) and Pangu Gu et al.
+(2023).
+which has been adapted for unanswerable questions, are the two SoTA models.
+For these, we use the available code.
+[^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in Appendix A.10.
+FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions.
+Instead of retrieve-then-generate, KB-Binder Li et al.
+(2023) follows a generate-then-ground approach.
+It is the SoTA for in-domain few-shot KBQA, and overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al.
+(2024).
+We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA since source code is not available.
+To adapt these two baselines for unanswerability, for fair comparison, we modify their logical form generation prompt in the same fashion as PUn for FUn-FuSIC.
+Additionally, for FuSIC-KBQA, we remove execution-guided feedback (EGF) since it fails for unanswerability.
+We denote this model FuSIC-KBQA-U.
+Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, without FUn.
+More details about use of KB-Binder are in the Appendix A.12.
+We use $\mathcal{L}=$gpt-4-0613 for all three models.
+For a fair comparison, we provide all three LLM-equipped models the same aggregated prompt limit for a question.
+To achieve this, to FUn-FuSIC we provide zero-shot generation and $n=4$ FUn iterations, to FuSIC-KBQA-U 5-shot generation, and to KB-Binder 25-shot generation.
+Though FUn-FuSIC and FuSIC-KBQA allow flexible use of multiple supervised retrievers, for meaningful comparison with RetinaQA, we adapt RetinaQA as retriever for FUn-FuSIC and FuSIC-KBQA-U.
+More details about KB-Binder and FuSIC-KBQA’s generation is in Appendix A.12, and details of FuSIC-KBQA’s retriever along compute infrastructure are in Sec.
+A.9.
 Evaluation Measures:
-Following existing literature Patidar et al.
+Since our primary task is generation of logical forms, our primary focus in evaluation is on logical forms as well.
+For logical form evaluation, the existing exact match measure (EM) Ye et al.
+(2022) works only for s-expression as the language.
+Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression, we propose a new measure that works for both languages.
+This measure, EM-s, considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answers upon execution.
+This is a necessary but not sufficient condition for logical form equivalence.
+So, we compared EM-s with EM in settings where both are applicable, and found $>98\%$ agreement.
+More details are in the Appendix A.3.
+Since logical form evaluation is not completely accurate (logical forms different from the gold one may still be correct), following existing literature Patidar et al.
 (2023); Faldu et al.
-(2024), we evaluate answers using regular and lenient F1 (F1(R) and F1(L)).
-F1(L) additionally gives credit for the original answer for the complete KB.
-For logical form evaluation, the existing evaluation measure (EM) Ye et al.
-(2022) works only for s-expression as the language.
-Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL and RetinaQA in s-expression, we propose a new measure (EM-s) that works for both.
-EM-s considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answer sets upon KB execution.
-Since this is a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable, and found $>98\%$ agreement.
-More details are in the Appendix (Sec. C).
+(2024), we also evaluate answers using regular F1 (F1(R)), also lenient F1 (F1(L)), which does not penalize the original answer for the complete KB.
 Answerable Unanswerable
@@
 FUn-FuSIC 57.0 46.0 92.0 90.0 62.0
-sc 63.0 52.0 66.0 64.0 36.0
+scUn $\Rightarrow$ sc 63.0 52.0 66.0 64.0 36.0
 w/o syntax 57.0 46.0 92.0 90.0 62.0
@@
 w/o ans-inc 55.0 44.0 92.0 90.0 62.0
-Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on WebQSP $\rightarrow$ GraphQAbility.
-sc denotes replacing self consistency for unanswerability with self consistency.
-Other rows remove checks for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
+Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on subset of WebQSP $\rightarrow$ GraphQAbility.
+scUn $\Rightarrow$ sc denotes replacing scUn with self consistency.
+Other rows remove verifiers for syntax error (w/o syntax) (V1), KB inconsistency (w/o kb-inc) (V2), question logical form disagreement (w/o q-lf) (V3) and answer incompatibility (w/o ans-inc) (V4).
 5.2 Unanswerability Setting
@@
 1.
-We make the following observations.
-(a) For answerable questions, FuSIC-U and FUn-FuSIC significantly outperform RetinaQA.
-Across the two datasets, FUn-FuSIC mostly performs better than FuSIC-U (both F1 and EM-s).
-(b) For unanswerable questions, FuSIC-U performs well in terms of the F1 scores, but not EM-s, indicating empty answers with incorrect logical forms.
-RetinaQA performs well in terms of F1 scores and EM-s.
-FUn-FuSIC performs slightly worse than RetinaQA for WebQSP $\rightarrow$ GrailQAbility, but significantly better for WebQSP $\rightarrow$ GraphQAbility.
-(c) Overall, FUn-FuSIC significantly outperforms both baselines in terms of EM-s, and is slightly below FuSIC-U in terms of F1(L) for WebQSP $\rightarrow$ GrailQAbility.
-This establishes usefulness of FUn for few-shot transfer KBQA with unanswerability.
+Note that the Overall columns determine superiority of one model over another, while the Answerable and Unanswerable columns provide further drill-down for analysis.
+First, we observe that FUn-FuSIC significantly outperforms all baselines in terms of EM-s, and performs at par with FuSIC-KBQA-U and significantly better than all other models in F1.
+However, the other LLM-equipped models do not far surpass supervised models.
+In fact, all 4 models perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility.
+This establishes usefulness of FUn+scUN for few-shot transfer KBQA with unanswerability.
+Secondly, each model trades off performance differently between answerable and unanswerable questions.
+RetinaQA, Pangu and also KB-Binder fare better for unanswerable questions, while FUn-FuSIC and FuSIC-KBQA-U fare better for answerable ones.
+However, FUn-FuSIC achieves the best balance between the two.
 We next briefly address research question R2.
@@
 2.
-While all models struggle with data-level unanswerability (EM-s), RetinaQA performs the worst and FuSIC-U is the best.
-For schema-level unanswerability, FUn-FuSIC is the best in one dataset, while RetinaQA is the best in the other.
-This indicates that more work is needed for robustness across unanswerability categories, and simultaneous use of PUn (prompting) and FUn (feedback) is the most promising approach.
+Models struggle to fare well across both data-level and schema-level unanswerability.
+FuSIC-KBQA-U performs the best for data-level while performing poorly (in terms of EM-s) for schema-level.
+Conversely, RetinaQA performs well for schema-level, but has poor data-level EM-s.
+FUn-FuSIC performs poorer for data-level compared to schema-level but achieves the best balance by far between the two categories among all models.
 We next address research question R3.
@@
 3 records performance for different ablations of FUn-FuSIC.
-The biggest benefit, for both answerable and unanswerable questions, comes from question logical form disagreement feedback.
-Feedback on KB inconsistency and answer incompatibility also make significant contributions to the performance.
+Note that this experiment is performed on a subset of the test data (50 questions drawn randomly from the each of the answerable and unanswerable categories).
+The biggest benefit, for both answerable and unanswerable questions, comes from Question - Logical Form Disagreement verifier.
+KB Inconsistency and Answer Incompatibility verifiers also make significant contributions to the performance.
+This demonstrates the usefulness of weak verifiers.
+Replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
 Model WebQSP $\rightarrow$ WebQSP $\rightarrow$
 GrailQA-Tech GraphQA-Pop
-FuSIC 70.8 52.3
+FuSIC-KBQA 70.8 52.3
 FUn-FuSIC(sc) 73.6 67.0
-FuSIC-U 62.6 43.4
+FuSIC-KBQA-U 62.6 43.4
 FUn-FuSIC(scUn) 71.2 65.0
 Table 4: Performance using F1 of different models for few-shot KBQA transfer with only answerable questions.
-The models in the first two rows assume knowledge of answerability of all questions, while the those in the next two do not.
+The models in the top block have prior knowledge of answerability, while those in the bottom block do not.
 5.3 Answerable Setting
-We now address answerable-only KBQA and research question R4.
-We use two of the datasets used in existing literature for this setting, including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop) Patidar et al.
-(2024) [^3] https://github.com/dair-iitd/FuSIC-KBQA/.
-For consistency with existing literature, here we use TIARA Shu et al.
-(2022) as the retriever for all models.
-This setting can be evaluated in two ways.
-In the first setting, the models are aware that all questions are answerable.
-This is the setting studied in the KBQA literature, and FuSIC-KBQA is the established SoTA model outperforming a host of supervised and LLM-based models adapted for the task Patidar et al.
-(2024).
-We appropriately adapt FUn-FuSIC as discussed earlier in Sec.
-3.2.
+We now address research question R4 (answerable-only KBQA).
+We use two datasets from existing literature Patidar et al.
+(2024), including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop).
+[^5] https://github.com/dair-iitd/FuSIC-KBQA/ For consistency with existing literature, here all models use TIARA Shu et al.
+(2022) as the retriever.
+This setting admits two sub-settings: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this prior knowledge.
+Setting (A) has been studied for KBQA Patidar et al.
+(2024), and FuSIC-KBQA is the established SoTA model, outperforming a host of supervised and LLM-based models adapted for the task.
+In this setting, FUn-FuSIC requires three simple modifications.
+(i) PUn is replaced with prompt for answerability, (ii) In FUn, category of V4b (empty answer) changes from weak verifier to strong verifier, and (iii) scUn is replaced by self-consistency.
 The first two rows in Tab.
-4 record performance for this setting.
-We see that FUn-FuSIC with self-consistency significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting.
-This establishes the usefulness of the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks.
-In the more practical second setting, which has not been studied before, although all questions are answerable, the models are not aware of this and make predictions assuming unanswerability.
-For this we evaluate FuSIC-U and FUn-FuSIC as in Sec.
-4.2.
+4 record performance for setting (A).
+FUn-FuSIC significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting.
+This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for KBQA transfer, even without unanswerability.
+In the realistic second setting (B) – not studied before – the models make predictions assuming unanswerability.
+Here, we evaluate FuSIC-KBQA-U and FUn-FuSIC as in Sec.
+5.2.
 The bottom two rows of Tab.
@@
 We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin.
-This further establishes the usefulness of scUn when guarantees about answerability do not exist.
+This further establishes the usefulness of scUn when guarantees about answerability are not available.
 5.4 Error Analysis
-We have performed detailed error analysis for FUn-FuSIC.
-For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions with EM-s $<1$ (46.2% of questions).
-There are two high-level error categories.
-(E1) Retrieval Error: At least one necessary KB element (class, relation, entity) is missing in the retrieved KB context (23.4% of questions).
-(E2) Generation Error: Despite correct retrieval, the generated logical form is incorrect (22.8% of questions).
-Generation errors can be further sub-categorized.
-(E2a) The gold logical form is NK, but some (non-NK) logical form is predicted (8.4% of questions).
-(E2b) A (non-NK) gold logical form exists, but NK is predicted, (4.6% of questions).
-(E2c) The gold and the predicted logical forms are non-NK, but different (9.8% of questions).
+EM-s $<1$ 46.2
+Retr. Err. 23.4
+Gen. Err. 22.8
+$l^{*}=$ NK, ${\hat{l}\neq}$ NK 8.4
+$l^{*}\neq$ NK, ${\hat{l}=}$ NK 4.6
+$l^{*}\neq$ NK, ${\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
+Table 5: FUn error analysis on WebQSP $\rightarrow$ GraphQAbility.
+$l^{*}$ ${\hat{l}}$ denote gold & generated logical forms.
+Retrieval error means retrieval $r$ is missing $\geq 1$ KB elements (class, relation, entity) necessary for $l^{*}$.
+Generation error implies ${\hat{l}}\neq l^{*}$ despite correct retrieval.
+For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions whose logical forms are judged incorrect (EM-s $<1$).
+See results in Tab.
+5.
 We found three main causes for generation errors.
-(1) Some questions are inherently ambiguous, admitting multiple valid logical forms in a complete KB.
-Incompleteness eliminates one of these, so that the gold logical form is NK, while the model identifies the other still existing non-NK logical form.
-(2) FUn is unable to generate the gold logical form within its maximum number of iterations ($n=4$).
-(3) FUn generates the gold logical form but it is rejected incorrectly by the question logical form disagreement checker.
+(1) Some questions are inherently ambiguous, admitting multiple valid logical forms $l_{1}$ and $l_{2}$ in the original complete KB, though only one is recognized as the gold ($l^{*}=l_{1}$).
+Deletion to introduce unanswerability eliminates $l_{1}$, so we update $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized.
+(2) $l^{*}=l_{1}$, ${\hat{l}}=l_{2}$, where $l_{1}\neq l_{2}$ while being semantically equivalent, but are incorrectly judged non-equivalent by EM-s.
+(3) FUn is genuinely unable to generate $l^{*}$ or any semantic equivalent within its iteration limit.
+We also estimated the accuracy of specific weak verifiers by manual analysis of $\sim$100 instances from test.
+Verifier V3 for Question - Logical Form Disagreement has accuracy of 93.6% overall, with its Query to Natural Language Back-translation component having 90%.
+Verifier V4b for Empty Answer has 67.8% (this is expected since $\sim$25of questions have ($l^{*}\neq$ NK, $A^{*}=$ NA).
 6 Conclusions
-For robust, low-resource KBQA, we have proposed the practical task of few-shot transfer learning with unanswerability.
-We have proposed FUn-FuSIC that adapts the existing SoTA model that assumes answerability by (i) candidate logical form generation by iterative LLM prompting with feedback from diverse error checks – including a novel back-translation based equivalence check, and (ii) assessing confidence in the candidates by adapting self consistency to detect answerability and different categories of unanswerability.
-Using newly created datasets, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable KBQA.
-Our error analysis suggests that performing well across all categories of unanswerability in the few-shot transfer setting is still a challenge and should be a focus of further research.
+For real-world robust, low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability.
+We have proposed FUn-FuSIC that augments for unanswerability the SoTA few-shot answerable-only KBQA transfer model by (i) iterative repair using feedback from a suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assessing this candidate set to detect unanswerability (and its category) and/or identify the best logical form using self consistency adapted for unanswerability.
+Using two newly created datasets for this novel task, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable few-shot transfer KBQA.
+Error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge and should be a focus of further research.
 7 Limitations
@@
 At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge.
-However, can incorrectly inferring unaswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk?
+However, can incorrectly inferring unanswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk?
 While we cannot imagine such a risk at the present time, this may require more thought.
@@
 We begin by defining the few-shot transfer learning setting for Knowledge Base Question Answering (KBQA) with unanswerability, and then present our solution for it.
-A.1.1 KBQA:
+A.1.1 KBQA: Detailed Background
 A Knowledge Base $G$ consists of a schema with data stored under it.
@@
 We are given a target Knowledge Base $G^{t}$ (consisting of entity types $T^{t}$, relations $R^{t}$, entities $E^{t}$ and facts $F^{t}$) and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns a non-empty answer $A^{t}$ for the question $q^{t}$.
-A.1.2 Few Shot Transfer Learning for KBQA:
+A.1.2 Few Shot Transfer Learning for KBQA
 In few-shot transfer learning, we are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms in the target domain.
@@
 Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
-A.1.3 KBQA with unanswerability:
+A.1.3 KBQA with Unanswerability
 A natural language question q is said to be answerable for a Knowledge Base $G^{t}$ if it has a corresponding logical form $l$ which when executed over $G^{t}$ returns a non-empty answer $A$.
@@
 Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer.
-Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Entity: The topic entity specified in the question is missing from the KB.
-Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediated entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB.
+Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Topic Entity: The topic entity specified in the question is missing from the KB.
+Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediary entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB.
 However, the (subject, relation, object) path is not connected in the KB.
-A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability:
-This setting is a generalisation of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$.
+A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability
+This setting is a generalization of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$.
 The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form.
 The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
-A.2 Algorithm
-Figure 2: Feedback Algorithm for Robustness in KBQA
-1:Given: KB-Engine, KB-Ontology, qn $\triangleright$ qn is natural language query
-[…]
-3:$L\leftarrow[]$
-4:for $i=0$ to $N$ do
-5:     $lf\leftarrow LLM(prompts)$
-6:     $syntax\_error\leftarrow KB-Engine(lf)$
-7:     if $syntax\_error\neq\text{null}$ then
-8:         $syntax\_prompt\leftarrow SynPrompt(lf,syntax\_error)$
-9:         $prompts.append(syntax\_prompt)$
-10:         continue
-11:     end if
-12:     $kb\_consistency\_error\leftarrow CheckTypes(lf,KB-Ontology)$
-13:     if $kb\_consistency\_error\neq\text{null}$ then
-14:         $kb\_consistency\_prompt\leftarrow LFPrompt(lf,kb\_consistency\_error)$
-15:         $prompts.append(kb\_consistency\_prompt)$
-16:         continue
-17:     end if
-18:     $qlf\_error\leftarrow CheckMeaning(lf,qn)$
-19:     $ans\_consistency\_error\leftarrow CheckAnswer(KB-Engine(lf),lf.nodes)$
-20:     if $qlf\_error\neq\text{null}$ then
-21:         $qlf\_prompt\leftarrow NLPrompt(lf,qlf\_error)$
-22:         $prompts.append(qlf\_prompt)$
-23:         if $ans\_consistency\_error$ = null then
-24:              $L.append(lf)$
-25:         end if
-26:         continue
-27:     end if
-28:     if $ans\_consistency\_error\neq\text{null}$ then
-29:         $ans\_consistency\_prompt\leftarrow EGF(lf,egf\_error)$
-30:         $prompts.append(ans\_consistency\_prompt)$
-[…]
-32:              $L.append(lf)$
-33:         end if
-34:         continue
-35:     end if
-36:     return $lf$triangleright all checks passed
-37:end for
-38$\triangleright$ despite N-tries, no lf passes
-39:$L_{ne}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)\neq\text{null}]$
-40:$L_{e}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)=\text{null}]$
-41:$poss\_answers\leftarrow[KB-Engine(lf)\text{ for }lf\text{ in }L_{ne}]$
-42:if $scUn(poss\_answers)\neq\text{null}$ then $\triangleright$ No Answer-aware-self-consistency
-43:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==scUn(poss\_answers)])$
-44:else if $is\_unanswerable$ then
-45:     if $L_{e}\neq[]$ then
-46:         return $SelectBestLF(qn,L_{e})$
-47:     end if
-48:     return NK
-49:else
-50:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==sc(poss\_answers)])$triangleright self-consistency, with tie breaking
-51:end if
-Algorithm 1 Self-Consistency Algorithm
-1:Given: $possible\_answers$
-2:Filter out null values from possible_answers:
-3:$possible\_answers\leftarrow[p\text{ for }p\text{ in }possible\_answers\text{ if }p\neq\text{null}]$
-4:if $possible\_answers=[]$ then
-5:     return null
-6:end if
-7:$MajVote\leftarrow MaxFreqItem(possible\_answers)$
-8:return $MajVote$
-Algorithm 2 Self-Consistency with Unanswerability Algorithm
-1:Given: $possible\_answers$
-2:$MajVote\leftarrow self-consistency(possible\_answers)$
-3:if $MajVote=\text{null}$ then
-4:     return null
-5:end if
-6:$count\leftarrow Cnt(MajVote,possible\_answers)$
-7:if $count>\frac{\text{len}(possible\_answers)}{2}$ then
-8:     return $MajVote$
-9:else
-10:     return null
-11:end if
-A.3 Working Example of The Algorithm
-Question: Where all were the Olympics held before 2024?
-Logical Form 1 ?y{ ?x olympics.venue ?y
-    ?y olympics.year ?z
-    FILTER(?z = 2024)}
-Feedback 1 Type: Logical Form Semantic Error Description: ?y is assigned types [’venue’, ’olympics’] by relations [’olympics.venue’, ’olympics.year’]. These types are mutually incompatible.
-Logical Form 2 ?y{ ?x olympics.venue ?y
-    ?x olympics.year ?z
-    FILTER(?z = 2024)}
-Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked returns all places where the Olympics were held before 2024.
-However, the question we answer returns the places where the Olympics were held in 2024.
-Logical Form 3 ?y{ ?x olympics.venue ?y
-    ?x olympics.year ?z
-    FILTER(?z < 2024)}
-Feedback 3 Status: Passes All Checks Details: lf3 satisfies syntax, logical form semantic, natural language semantic and execution checks.
-Appendix B Ablations
-We perform ablations of our best performing model - FUn-FuSIC by removing different types of feedback from the model.
-We report performance on 100 questions(50 answerable + 50 unanswerable) sampled from the WebQSP $\rightarrow$ GraphQAbility dataset.
-Refer to 5
-Ablation WebQSP $\rightarrow$ GraphQAbility
-Overall Ans Unans
-F1(L) SM F1 SM F1(L) F1(R) SM
-FUn-FuSIC 74.5 54.0 57.0 46.0 92.0 90.0 62.0
-sc 64.5 44 63.0 52.0 66.0 64.0 36.0
-w/o syntax 74.0 54.0 57.0 46.0 92.0 90.0 62.0
-w/o kb-inc 71.8 43.0 51.7 42.0 92.0 88.0 44.0
-w/o qf-lf 51.7 24.0 47.7 38.0 55.8 48.0 10.0
-w/o ans-inc 73.5 53.0 55.0 44.0 92.0 90.0 62.0
-Table 5: Detailed ablation Performance of FUn-FuSIC on WebQSP $\rightarrow$ GraphQAbility.
-sc denotes replacing self consistency for unaswerability with standard self consistency, while other rows indicate removing check for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
-Appendix C EM-s: EM for sparql
+A.2 Error Checks
+(V2) Semantic Error (KB Inconsistency):
+A syntactically correct logical form $l$ may still be inconsistent with the schema of $G$.
+This is a likely error even for SoTA LLMs since these are unfamiliar with the specific KB $G$.
+Semantic errors have different categories, such as type-incompatibility, schema hallucinations.
+(V2a) Incompatibility in types: $l$ contains a variable and a connecting relation whose types are incompatible in $G$.
+This is the case for LF1 for all three KBs in the example.
+(V2b) Schema hallucinations: $l$ contains schema elements (types, relations, entities) absent in $G$.
+(V2c) Type casting errors: Literals in $l$ are not correctly type cast for $G$, e.g.
+numeric literals as float for Freebase.
+All of these are certain checks, and are implemented using rules defined over $l$ and $G$.
+The feedback mentions the type of error and the specifics, e.g., the hallucinated relation, or the incompatible type-relation pair.
+(V3) Question Logical form Disagreement:
+FUn performs equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline.
+(i) The variable names in $l$ are first naturalized to $l^{n}$ considering $q$ and preserving semantics, e.g.
+by replacing ’?x’ with ’?actor’.
+(ii) $l^{n}$ is back-translated into a natural language question $q^{b}$.
+(iii) $q^{b}$ is finally checked for semantic equivalence with $q$.
+The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$.
+More details about few-shot construction are in Sec.
+A.8.2.
+The feedback mentions lack of equivalence as the type of error.
+(V4) Answer Inconsistency:
+If the $l$ is syntactically and semantically correct, it is executed over $G$ to obtain an answer $a$.
+$a$ is then checked for compatibility with $q$.
+This may fail for different reasons.
+(V4a) $a$ (which is a set in general) contains an entity also in $l$ and therefore mentioned $q$, which is an aberration.
+(V4b) $a$ is empty, as in LF3 for KB2 in the example.
+All of these checks are implemented using rules defined over $l$ and $G$.
+Note that while the first two are certain checks, the last is not.
+An empty answer is valid for unanswerable questions, as for LF3 for KB2, but invalid for answerable ones.
+As before, the feedback mentions the type of error and the specifics.
+A.3 EM-s: Automated check for equivalence of SPARQL programs
 As has been observed in Patidar et al.
@@
 Since these are a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable and found $>98\%$ agreement.
-Appendix D Performance across different categories of unanswerability
+A.4 Performance across different categories of unanswerability
 We dive deeper into the different categories of unanswerability, as explained in Patidar et al.
@@
 Based on performance on the WebQSP $\rightarrow$ GrailQAbility and WebQSP $\rightarrow$ GraphQAbility datasets, we validate the that the trends are indeed as expected.
-Appendix E Cost Analysis
+A.5 Cost Analysis
 FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as U-FuSIC and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16.
@@
 Hence, the two models are comparable in terms of cost.
-Appendix F FuSIC-KBQA: High Level Overview
+A.6 FuSIC-KBQA Details
 Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA.
-FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (b) LLM ReRanking: an LLM is used to select the top-10 classes, top-10 relations and top-5 paths among the top-100 elements as ranked by the supervised retriever, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
-Appendix G Prompts
-G.1 PUn prompt
+FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
+A.7 FUn-FuSIC Prompts
+Here we provide details of various prompts used by FUn-FuSIC.
+A.8 PUn prompt
+The first prompt is for Prompting for Unanswerability (PUn).
 Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively.
@@
 3.
-When you will generate a new sparql only using the candidate entities, relations and entity types.For entity type check please use this relation "type.object.type".Do not use entity names in the query.
+When you will generate a new sparql only using the candidate entities, relations and entity types.
+For entity type check please use this relation "type.object.type".D o not use entity names in the query.
 Use specified mids.
-If it is impossible to construct a query using the provided candidate relations or types, return "NK"Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
+If it is impossible to construct a query using the provided candidate relations or types, return "NK".
+Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
 NK exemplar Question: the tv episode segments spam fall under what subject?
@@
 sparql:
-G.2 FUN prompt
-syntax error feedback Correct the syntax of the following sparql query. Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ... Virtuoso error: word AND not defined
-Target KB Disagreement feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query.
+A.8.1 FUN prompt
+Syntax error(V1) Feedback Correct the syntax of the following sparql query.
+Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ...
+Virtuoso error: word AND not defined
+KB Inconsistency(V2) Feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query.
 The assigned relation types by [’computer.computer_emulator.computer’, ’type.object.type computer.computer_peripheral’] are [’computer.computer’, ’computer.computer_peripheral’].
@@
 DO NOT APOLOGIZE - just return the best you can try.
-Question Logical form disagreement feedback The question that you answer is NOT same as what you’ve been asked for!
+Question Logical form disagreement(V3) feedback The question that you answer is NOT same as what you’ve been asked for!
 You have answered the question "Which opera productions has Gino Marinuzzi conducted?" but you were asked to answer "what is the name of the premiere opera production conducted by gino marinuzzi?".
@@
 The next three prompts fall under the answer incompatibility feedback
-empty answer feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
-intermediate node feedback The generated sparql returns an intermediate type node when executed on the freebase KG.
+Answer Inconsistency(V4b) feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
+Intermediate Node(V4a) feedback The generated sparql returns an intermediate type node when executed on the freebase KG.
 Maybe the answer node is an adjacent node to what we currently query for.
 Please generate again a different executable sparql using the same context and constraints.
-answer node in question feedback The logical form upon execution returns International System of Units, which is not answering the question.
+Answer Inconsistency(V4a) feedback The logical form upon execution returns International System of Units, which is not answering the question.
 Please reconstruct the query using same context and constraints.
-G.3 Question Logical Form Agreement Prompt
+A.8.2 Prompt for Verifier V3 Question Logical Form Agreement
 The few shots provided for verifying question logical form agreement come from the few shots provided in the target domain $D^{t}$.
@@
 Firstly, we perform back-translation to obtain natural language question from the logical form
-Naturalization of variable names change the sparql query to have variable names representative of what objects they refer to.
+Naturalization of variable names(V3(i)) change the sparql query to have variable names representative of what objects they refer to.
 transform the variable names in this query.
@@
 Some examples of few shots are provided below-
-question logical form agreement check Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals?
+Question Logical Form Sgreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals?
 Question originally asked: name the cricket player who has 31 odi stumps.
@@
 explanation:
-G.4 Logical Form Selection Prompt
+A.8.3 Prompt in scUn for Logical Form Selection
 We store the back-translated natural language questions along with each predicted logical form.
-Here, we prompt the LLM to select the closest back-translated natural language question to the orginal question.
+Here, we prompt the LLM to select the closest back-translated natural language question to the original question.
 Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director?
@@
 Even if none is very close, return the one that is semantically closest?
-pls explain your answer as well
-Appendix H Supervised Models training details
+Please explain your answer as well
+A.9 Supervised Models training details
 We use Hugging Face  Wolf et al.
 (2020), PyTorch Paszke et al.
-(2019) for our experiments and use the Freebase setup specified on github 555 https://github.com/dki-lab/Freebase-Setup .
+(2019) for our experiments and use the Freebase setup specified on github 777 https://github.com/dki-lab/Freebase-Setup .
 We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM.
@@
 (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^10] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^11] https://github.com/dair-iitd/RetinaQA.
-Appendix I Supervised Models inference details
+A.10 Supervised Models inference details
+We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping.
+In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability.
+We again use the target dev sets for this.
 We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths.
@@
 (b) determining the threshold value.
-RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labelled as NK.
-We choose the optimal value of the threshold to maximise the overall EM-s score over the dev set.
+RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK.
+We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
+A.11 Pangu Adaptation Details
+Similar to RetinaQA, We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping.
+We use 1 GPU for training.
+Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability.
+Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK.
+We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
+A.12 KB-Binder and FuSIC-KBQA Adaptation Details
+For KB-Binder, we make use of publicly available code  [^12] https://github.com/ltl3A87/KB-BINDER (MIT License).
+We use self-consistency and majority voting with 6 examples, as in the experiments in the paper.
+In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data.
+We restrict its retrieval to our target training set $D_{t}$ with 25 examples.
+KB-Binder reports experiments using code-davinci-002 as the LLM.
+For consistency and fair comparison, we replace these with gpt-4-0613 as in our approach.
+This model generates logical forms in s-expression, which we preserve.
+For FuSIC-KBQA Patidar et al.
+(2024), since no code is available for this model, we use our own implementation based on the description in the paper.
+For FuSIC-KBQA, and FUn-FuSIC we use $temperature=0$.
+A.13 Real Example of FUn
+Question: Where all were the Olympics held before 2024?
+Logical Form 1 ?y{ ?x olympics.venue ?y
+    ?y olympics.year ?z
+    FILTER(?z = 2024)}
+Feedback 1 Type: Logical Form Semantic Error Description: ?y is assigned types [’venue’, ’olympics’] by relations [’olympics.venue’, ’olympics.year’]. These types are mutually incompatible.
+Logical Form 2 ?y{ ?x olympics.venue ?y
+    ?x olympics.year ?z
+    FILTER(?z = 2024)}
+Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked returns all places where the Olympics were held before 2024.
+However, the question we answer returns the places where the Olympics were held in 2024.
+Logical Form 3 ?y{ ?x olympics.venue ?y
+    ?x olympics.year ?z
+    FILTER(?z < 2024)}
+Feedback 3 Status: Passes All Checks Details: lf3 satisfies syntax, logical form semantic, natural language semantic and execution checks.
+A.14 Algorithm for Self Consistency with Unanswerability
+The high level algorithm for self consistency with Unanswerability (scUn) is described in Algo. 3.
+Algorithm 3 ScUn$(q,L,t,\mathcal{L})$
+1:$(c,l,A)=\mbox{assessConf}(q,L,t,\mathcal{L})$
+2:if ($\mbox{{\bf if} }(c)$, $\mbox{{\bf return}}(l,A)$
+[…]
+The high level algorithm for assessing confidence in the set of candidate logical forms (assessConf) is described in Algo. 4.
+Algorithm 4 $\mbox{assessConf}(q,L,t,\mathcal{L})$
+1:$(c,L^{p},A^{p})=\mbox{popAnsNE}(L,t)$
+2:if $(c)$ then
+3:     $l=\mbox{selectBestNE}(q,L^{p},\mathcal{L})$
+4:     $\mbox{{\bf return}}(\mbox{True},l,A^{p})$
+5:end if
+6:$(c,L^{p})=\mbox{popAnsE}(L,t)$
+7:if $(c)$ then
+8:     $l=\mbox{selectBestE}(q,L^{p},\mathcal{L})$
+9:     $\mbox{{\bf return}}(\mbox{True},l,\texttt{NA})$
+10:end if
+[…]

### v3

@@
 Abstract
-Real-world applications of KBQA require models to handle unanswerable questions with a limited volume of in-domain labeled training data.
-We propose the novel task of few-shot transfer for KBQA with unanswerable questions and contribute two new datasets for performance evaluation.
-We present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA, the state-of-the-art few-shot transfer model for answerable-only KBQA.
-We first note that FuSIC-KBQA’s iterative repair makes a strong assumption that all questions are answerable.
-As a remedy, we propose Feedback for Unanswerability (FUn), which uses iterative repair using feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerabilty to better assess the answerability of a question.
-Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA for answerable few-shot transfer as well.
+Real-world applications of KBQA require models to detect different types of unanswerable questions with a limited volume of in-domain labeled training data.
+We propose the novel task of few-shot transfer for KBQA with unanswerable questions.
+The state-of-the-art KBQA few-shot transfer model (FuSIC-KBQA) uses an iterative repair strategy that assumes that all questions are answerable.
+As a remedy, we present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA with Feedback for Unanswerability (FUn), which is an iterative repair strategy for answerable as well as unanswerable questions.
+FUn uses feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerability for assessing answerability of questions.
+Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA performance for answerable few-shot transfer as well.
+We have made datasets and other resources publicly available [^1] https://github.com/dair-iitd/FUn-FuSIC
 Iterative Repair with Weak Verifiers
 for Few-shot Transfer in KBQA with Unanswerability
-Riya Sawhney, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
+Riya Sawhney‡,§, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, §Graviton Research Capital, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
 1 Introduction
 The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes as input a Knowledge Base (KB) and a natural language question, and outputs a logical form (or program) that produces the answer upon execution over the KB.
-KBQA has important real-world applications, which can require systems to be low resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically the ability to identify questions that cannot be answered based on existing KB.
+KBQA has important real-world applications, which require KBQA systems to be low-resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically able to identify questions that cannot be answered from the KB.
 Traditional supervised models (e.g., Ye et al.
@@
 (2023); Nie et al.
-(2024) fall short in both aspects.
+(2024) for KBQA fall short in both aspects.
 Limited recent work has addressed these independently – in-domain methods for KBQA with unanswerability trained with large labeled data Patidar et al.
 (2023); Faldu et al.
-(2024), and FuSIC-KBQA for few-shot transfer assuming questions are answerable Patidar et al.
+(2024), and FuSIC-KBQA for few-shot transfer assuming answerable questions Patidar et al.
 (2024).
-No existing work simultaneously addresses both desiderata in a single KBQA system.
+No existing single KBQA model simultaneously addresses both desiderata.
 In response, we propose the novel task of few-shot transfer learning for KBQA with unanswerability.
-Specifically, the target domain may offer only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
-For KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative error guided repair.
-Specifically, multiple iterations of execution-guided feedback are run until a non-empty answer is obtained – this assumes that all questions are answerable.
-For our task, we first design FuSIC-KBQA-U, which, to address Unanswerability, eschews the now-inappropriate repair step, and modifies the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars.
+Specifically, the target domain has only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
+For few-shot KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative execution-error-guided repair.
+Specifically, multiple feedback-guided repair iterations are executed, checking emptiness of answers obtained by executing the generated program as indication of correctness, until a non-empty answer is obtained.
+This naturally fails when questions are allowed to be unanswerable.
+A simple-fix for addressing Unanswerability (FuSIC-KBQA-U) is to drop the inappropriate repair step, and modify the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars.
 Unlike in studies for unanswerability in general QA Slobodkin et al.
-(2023), this vanilla adaptation mostly generates incorrect logical forms for unanswerable questions.
+(2023), we found that FuSIC-KBQA-U mostly generates incorrect logical forms for unanswerable questions.
 As a remedy, we design a novel solution: FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA).
@@
 For logical forms, we use strong verifiers to identify obvious syntactic and semantic errors.
-We also propose a novel verifier: non-equivalence of the original question and the back-translation of the logical form.
-This is weak, due to potential errors in back-translation as well as in equivalence classification.
-Using such iterative weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms.
-Simply selecting from this set the logical form with majority answer via self-consistency Wang et al.
+We also propose a novel verifier involving a 3-component LLM-based pipeline: non-equivalence of the original question and the back-translation of the logical form.
+This verifier is also weak, due to potential errors in back-translation as well as in equivalence classification.
+Using such iterative strong and weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms.
+For selecting the consensus logical form from this set, using the majority answer as in self-consistency Wang et al.
 (2023) breaks down in the face of unanswerability.
-We introduce self-consistency for unaswerability, which additionally checks the likelihood of the majority answer, empty or otherwise, to produce the final output.
-For evaluation in our novel task, we create two KBQA transfer datasets with unanswerable questions in the target.
-Our experiments show that FUn-FuSIC comprehensively outperforms different types of SoTA models suitably adapted for this task, including LLM-based and more traditional models.
-We further find that iterative repair of logical forms using weak verifiers holds promise for KBQA in general.
-Using experiments over benchmark datasets for answerable-only few-shot KBQA transfer, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the known SoTA on the task.
+We introduce self-consistency for unanswerability, which assesses the likelihood of the majority answer, empty or otherwise, to select the consensus logical form.
+Since no datasets exist for our novel task, we create two new datasets for KBQA transfer with unanswerable questions.
+Our experiments show that FUn-FuSIC comprehensively outperforms different categories of SoTA models suitably adapted for this task, including LLM-based and more traditional models.
+We further find that iterative repair of logical forms using weak verifiers holds promise for even for KBQA with only answerable questions.
+Using experiments over benchmark datasets for this task, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the SoTA model for the task.
 In summary, our specific contributions are as follows.
 (a) We propose the problem of few-shot transfer for KBQA with unanswerability.
-(b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of weak verifiers.
+(b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of strong and weak verifiers.
 (c) We create new datasets for the proposed task, which we make public.
-[^1] https://anonymous.4open.science/r/FUn-FuSIC-1704 (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting.
-(e) We also show that for standard answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
+(d) We show that FUn-FuSIC outperforms adaptations of SoTA KBQA models for this new task.
+(e) We also show that even for answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
 2 Related Work
@@
 RetinaQA Faldu et al.
-(2024) is the SoTA model for KBQA unaswerability.
+(2024) is the SoTA model for KBQA unanswerability.
 However, this also requires large volumes of training data.
@@
 (2022), low-resource was originally not a focus.
-More recently, few-shot transfer for KBQA has been addressed Patidar et al.
+More recently, few-shot transfer for KBQA has been addressed by FuSIC-KBQA Patidar et al.
 (2024).
-Its FuSIC-KBQA model uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair.
-However, this formulation assumes answerability of all questions, as described in Section 1.
+FuSIC-KBQA uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair.
+However, this formulation assumes answerability of all questions.
 Simple LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al.
@@
 (2023b) use execution based refinement for program generation but without any notion of unanswerability.
-The extensive recent work on program self-repair using LLMs Olausson et al.
-(2024) also diverges strongly from our few-shot goal.
-It assumes availability of unit tests for every test instance.
-More importantly, as in the answerable KBQA literature, it assumes that a valid program exists for each specification.
-To the best of our knowledge, ours is the first work to apply iterative weak repair for KBQA, and to use iterative repair for unanswerability.
+FUn’s iterative repair idea may be useful in other natural language to program generation tasks where non-existence of a program with the required specification has not been studied to the best of our knowledge, such as NL-to-SQL  Dong et al.
+(2023); Pourreza and Rafiei (2023) and program self-repair using LLMs Olausson et al.
+(2024); Madaan et al.
+(2023); Grattafiori et al.
+(2024).
 3 Background & Problem Definition
-Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question $Q$ when executed over different KBs with $\leq 4$ iterations.
-$Q$ is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness).
-FUn iterations are shown using gray blocks, with Strong Verifiers in red and Weak Verifiers in green.
-(V1: Syntax Error omitted for brevity) scUn is shown using the blue block.
+Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question when executed over different KBs with 4$\leq$ 4≤ 4 iterations.
+The question is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness).
+In the KB depictions, the top graph represents the schema with different node colors for different entity types, and the bottom graph represents the data.
+(Names of different (real) books related to the author entity in the question are abbreviated in the data graph.) FUn iterations are shown in the gray blocks, with Strong Verifiers named using red text and Weak Verifiers using green text and $i$ denotes the iteration number.
+The outcome of verification is denoted as True or False, and the non-empty answer is shown when V4b returns False.
+The Syntax Error Verifier (V1) is omitted for brevity.
+scUn is shown in the blue blocks.
+The candidate logical forms ($L$) for sCun are shown at the top.
+For Non-empty Ans and Empty Ans agreement checks for, the outcome is either ‘Consensus’ or ‘No consensus’.
+At the end, FUn-FuSIC ‘Return’s a logical form (possibly NK) and an answer (possibly NA).
 A Knowledge Base (KB) $G$ consists of a schema and data.
@@
 The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$.
-Given a target KB $G^{t}$ and a natural language question $q^{t}$, the goal is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ (which is a set in general).
-Target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms.
-A related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
+Given a target KB $G^{t}$ and a natural language question $q^{t}$, the basic KBQA task is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ ($A^{t}\subset E$ in general).
+Other than SPARQL  Patidar et al.
+(2024), niche languages such as s-expressions  Li et al.
+(2023); Gu et al.
+(2023) are commonly used for logical forms in KBQA.
+In supervised in-domain KBQA, the target has large volumes of labeled training examples of questions and associated logical forms.
+For few-shot in-domain KBQA in contrast, target few-shots $D^{t}$ contain tens of labeled training examples.
+In few-shot transfer learning for KBQA Patidar et al.
+(2024), a related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
 Following Patidar et al.
-(2023), a question $q$ is answerable for a KB $G$ if it has a corresponding logical form $l$ which when executed over $G$ returns a non-empty answer $A$.
-A question is unanswerable, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$ (data-level unanswerability).
+(2023), a question $q$ is answerable for a KB $G$ if it admits a corresponding logical form $l$ which when executed over $G$ returns the ideal non-empty answer $A$.
+A question is unanswerable if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$, different from the ideal non-empty answer (data-level unanswerability).
 Schema-level unanswerability arises due to missing types and relations, while missing entities and facts lead to data-level unanswerability.
-More details are in Appendix  A.1.3.
-In KBQA with unanswerability, given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
-We now define few-shot transfer learning for KBQA with unanswerability.
+However, absence in $G$ of any entity mentioned in the question is categorized as schema level unanswerability, since it invalidates the logical form.
+More details are in Appendix  A.1.2.
+In KBQA with unanswerability Patidar et al.
+(2023); Faldu et al.
+(2024), given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
+In the supervised in-domain setting, this task involves large volumes of labeled training questions, containing both answerable and unanswerable, for the target.
+We now define our problem of interest: few-shot transfer learning for KBQA with unanswerability.
 A target question $q^{t}$ may be answerable or unanswerable due to missing schema or data in the target KB $G^{t}$.
-Target few shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories.
-However, most KBQA datasets contain only answerable questions, we model the source training data $D^{s}$ as containing only answerable questions.
-Now the unanswerability gap between the source and the target also needs to be bridged.
-More details are in Appendix A.1.
+Target few-shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories.
+The source training data $D^{s}$ has large volumes of labeled training data.
+Considering real world constraints, where most KBQA datasets contain only answerable questions, we assume that $D^{s}$ contains only answerable questions.
+Compared to the earlier few-shot KBQA transfer task definition, now there is additionally an unanswerability mismatch between the source and the target distribution.
+More details are in the Appendix (Sec.
+A.1).
 4 Proposed Approach: FUn-FuSIC
-Our proposed architecture FUn-FuSIC adapts FuSIC-KBQA for unaswerability.
+Our proposed model FUn-FuSIC preserves the basic architecture of FuSIC-KBQA and adapts its iterative repair strategy for unanswerability.
 The high-level algorithm is described in Algo.
 1.
-Preserving the retrieve-then-generate framework, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers.
-Each retriever $R_{i}$ is source-trained and then target fine-tuned.
+(Since the algorithms are not specific to the transfer task, we use $q$ instead of $q^{t}$ for brevity.) Preserving the retrieve-then-generate framework of FuSIC-KBQA, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers.
+Each retriever $R_{i}$ is source-trained and further target fine-tuned if required.
 The retrieval output $r$ of each $R_{i}$ consists of relevant schema elements (types and relations) for $q^{t}$, and data paths emanating from mentioned entities in $q^{t}$.
-The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few shots $D^{t}$.
+The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few-shots $D^{t}$.
 More details of the retrieval stage are in the Appendix (Sec.
-A.6).
-Algorithm 1 FUn-FuSIC$(q,G^{t},D^{t},R,V^{s},V^{w},\mathcal{L})$
-1:$r=\{\}$
+A.8.1).
 […]
-3:$l=\mbox{PUn}(\mathcal{L},I,q,r,D^{t})$
-4:$(e,l,A,L)=\mbox{FUn}(\mathcal{L},q,l,n,V^{s},V^{w},G^{t})$
-5:if ($\mbox{{\bf if} }(e)\mbox{ {\bf return}}(l^{*},A^{*})$
-6:else $\mbox{{\bf else} {\bf return}}(\mbox{scUn}(q,L,t,\mathcal{L}))$
-FUn-FuSIC makes three modifications to the FuSIC-KBQA architecture.
-First, the LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions.
+[…]
+[…]
+[…]
+[…]
+[…]
+6:else return $\mbox{scUn}(q,L,\mathcal{L})$
+FUn-FuSIC differs from FuSIC-KBQA in its iterative repair strategy in lines 3, 4 and 6 of Algo.
+1.
+The LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions.
 However, this simple approach is error-prone.
@@
 The detailed prompt $I$ is in the Appendix (Sec.
-A.8).
+A.11.1).
 We now come to the more significant modifications.
-First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unaswerability.
-This iterative repair, which we name Feedback for Unaswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed.
-For this, we introduce self-consistency for Unaswerability (scUn).
-scUn checks the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output.
+First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unanswerability.
+This iterative repair, which we name Feedback for Unanswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed for a consensus logical form and answer.
+For this, we introduce self-consistency for Unanswerability (scUn) (line 6).
+scUn assesses the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output.
 In the rest of this section, we describe FUn and then scUn.
@@
 A real example of FUn execution is in Sec.
-A.13.
+A.12.
 Algorithm 2 $\mbox{FUn}(q,l,n,V^{s},V^{w},G,\mathcal{L})$
-1:$i=0$, $k=0$, $L=\{\}$, $F=\mbox{""}$"
-2:while $++i\leq n$ do
+[…]
+2:while $i\leq n$ do
 3:  for $j=1\mbox{ to }k_{1}$ do
-4:         $(e,f)=V^{s}_{j}(l,q,G)$
+4:   $(e,f)=V^{s}_{j}(l^{(i)},q,G)$
 5:   $F=\mbox{Append}(F,f)$
@@
 8:  for $j=1\mbox{ to }k_{2}$ do
-9:         $(e,f)=V^{w}_{j}(l,q,G)$
+9:   $(e,f)=V^{w}_{j}(l^{(i)},q,G)$
 10:   $F=\mbox{Append}(F,f)$
-11:         if ($\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l\}$; $k++$
+11:   $\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l^{(i)}\}$; $k++$
 12:  end for
-13:     $l=Gen(\mathcal{L},I,q,F)$
-14:end while
-15:if ($\mbox{{\bf if} }(k=k_{2})\mbox{ {\bf return}}(\mbox{True},l,\mbox{Exec}(l,G),L)$
-16:else $\mbox{{\bf else} {\bf return}}(\mbox{False},l,\{\},L)$
+13:  $i=i+1$
+14:  $l^{(i)}=Gen(\mathcal{L},I,q,F)$
+[…]
+16:end while
+[…]
 The FUn algorithm is described in Algo.
 2 Starting with the initial logical form $l^{(0)}$, FUn performs at most $n$ verify-and-repair iterations to create a candidate set $L$ of probable logical forms.
-The example shows 3 FUn iterations for KB1 and KB2, and 2 iterations for KB3.
+Fig.
+1 shows 3 FUn iterations for KB1 and KB2, and 2 for KB3.
 In the $i^{th}$ iteration, FUn generates a new logical form $l^{(i)}$ by prompting $\mathcal{L}$ using $q^{t}$ and feedback $F$ received from checks in all previous iterations (line 13).
@@
 The strong verifiers $V^{s}$ are guaranteed to be correct, while the weak verifiers $V^{w}$ are potentially erroneous.
+$k_{1}$ and $k_{2}$ denote the total number of strong and weak verifiers respectively.
 The specific verifiers that we use in this paper are defined later in the section.
@@
 In the example, this happens for all three KBs in iteration 1.
-If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$.
+If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$ (denoted $\mbox{X}(l^{(i)},G)$).
 This happens in iteration 3 for KB3.
@@
 Logical form verifiers:
-FUn uses a suite verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$).
+FUn uses a suite of verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$).
 These may be syntactic, semantic, or execution-based, defined using simple rules or complex LLM functions over $l$, $q$ and $G$.
@@
 Additional details about the verifiers are in the Appendix (Sec.
-A.8.1 and Sec.
-A.2).
-Note however that FUn is a framework that is equally capable of working with a wholly different suite of verifiers.
+A.11.2 and Sec.
+A.3).
+Note that FUn is a framework that is capable of working with a wholly different suite of meaningful verifiers.
 (V1) Syntax Error:
-As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check: a valid logical form cannot have syntax error.
+As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check — a valid logical form cannot have syntax error.
 (V2) KB Inconsistency:
@@
 This verifier checks if $l$ is semantically equivalent to the original natural language question $q$.
-In our example, LF3 for KB1 disagrees with Q.
+In Fig.
+1, LF3 for KB1 disagrees with $q$.
 This is a weak verifier.
-First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities.
+First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities even when it is answerable.
 For example, $q$ mentions a person from a country, when $G$ has the relations born in and works in between these types.
-Secondly, this is implemented as a probabilistic classifier that can err.
-We define equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$.
-More details are in Sec.A.2.
+Secondly, this verifier is a probabilistic classifier that naturally makes occasional mistakes.
+We define equivalence check between $l$ and $q$ using a multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$.
+More details are in the Appendix (Sec.A.3).
 (V4) Answer Inconsistency:
@@
 Unless some logical form passes all checks and is therefore returned (Algo.
-2 line 6), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so.
+2 line 15), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so.
 For our specific suite of weak verifiers, $l^{(i)}$ is added to $L$ if it passes one of V4b ($A$ is non-empty) as for LF3 for KB1, or V3 ($l^{i}$ is equivalent to $q$), as for LF3 for KB2.
-Self Consistency for Unanswerability:
-If FUn fails to generate a single confident logical form at the end of $n$ iterations, as for KB1 and KB2, we use scUn to decide if the best candidate $l^{*}\in L$ has sufficient confidence.
-If scUn detects sufficient confidence, as for KB2, it outputs ($l=l^{*}$, $A=A^{*}$), where $A^{*}$ is the answer obtained upon executing $l^{*}$ and may be NA.
-If on the other hand, scUn decides insufficient confidence, as for KB1, it outputs ($l=$ NK, $A=$ NA).
-For identifying the consensus choice from $L$, one possibility is self-consistency Wang et al.
+Self Consistency for Unanswerability (scUn):
+Given a candidate set $L$ of logical forms and a question $q$, scUn assesses if the best candidate $l^{*}\in L$ has sufficient confidence.
+If so, it outputs ($l=l^{*}$, $A=A^{*}$), as for KB2, $A^{*}$ being the answer from executing $l^{*}$ (may be NA).
+Otherwise, scUn outputs ($l=$ NK, $A=$ NA), as for KB1.
+For identifying the consensus choice from $L$, one possibility is self-consistency (sc) Wang et al.
 (2023); Chen et al.
-(2023a) i.e., obtaining the answer for each $l\in L$, and returning those with the most common answer.
-Self-consistency requires some answer to accumulate enough probability by aggregation over reasoning paths.
-However, for unanswerable questions, no single answer accumulates sufficient probability, and self-consistency returns some low probability answer.
-To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$).
+(2023a) that considers the answer for each $l\in L$, and returns those with the most common answer.
+This requires some answer to accumulate enough probability by aggregation over reasoning paths.
+However, for unanswerable questions, no single answer accumulates sufficient probability, and sc returns some low probability answer.
+To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using a threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$).
 If so, scUn uses LLM prompting to select the most appropriate supporting logical form $l^{*}\in L$ considering $q$, and outputs ($l=l^{*}$, $A=A^{*}$).
-However, for KB1 in the example, the 3 logical forms among the candidates have 3 different answers, and therefore ($\lfloor\frac{|L|}{2}\rfloor=1$) no consensus emerges for non-empty answers.
-Then scUn considers logical forms from $L$ that agree on $A=$ NA.
+However, for KB1, the 3 logical forms among the candidates have 3 different answers, and therefore no consensus emerges ($\lfloor\frac{|L|}{2}\rfloor=1$).
+Here, scUn considers logical forms from $L$ that agree on $A=$ NA.
 If there are multiple such candidates, scUn selects the most suitable candidate $l^{*}$, again using LLM prompting, and outputs ($l=l^{*}$, $A=$ NA).
@@
 Further details on scUn are in the Appendix (Sec.
-A.14).
-.
+A.2).
 5 Experiments
-Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
-Overall Answerable Unanswerable Overall Answerable Unanswerable
+WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
+Model Overall Answerable Unanswerable Overall Answerable Unanswerable
 F1 EM-s F1 EM-s F1(L) F1(R) EM-s F1 EM-s F1 EM-s F1(L) F1(R) EM-s
 RetinaQA 58.4 42.2 28.7 26.0 88.0 84.8 58.4 49.7 35.8 18.7 15.2 80.7 78.7 56.4
-Pangu 54.5 43.8 31.23 29.6 83.8 80.4 58.0 53.4 33.0 30.3 26.4 76.5 74.8 39.6
+Pangu 54.5 43.8 31.2 29.6 83.8 80.4 58.0 53.4 33.0 30.3 26.4 76.5 74.8 39.6
 FuSIC-KBQA-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
@@
 Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability.
+Answerable and Unanswerable record performance for corresponding subsets and Overall for the entire dataset.
 We now present experimental evaluation of FUn-FuSIC.
-For few-shot KBQA transfer with unanswerability, we address the following research questions.
-(R1) How does FUn-FuSIC compare against SoTA models for KBQA, suitably adapted for this setting?
+First, for few-shot KBQA transfer with unanswerability, we address the following research questions.
+(R1) How does FUn-FuSIC compare against SoTA KBQA models suitably adapted for this setting?
 (R2) How does FUn-FuSIC perform across different categories of unanswerability?
 (R3) How do the different components of FUn-FuSIC contribute to its performance?
-In addition, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
+Then, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
 5.1 Experimental Setup
 Datasets:
-For few-shot KBQA transfer, available datasets have only answerable questions Patidar et al.
-(2024), while our task needs the target dataset to contain unanswerable questions as well.
-So we construct our own transfer datasets, augmenting existing ones.
+For in-domain and answerable KBQA, the three most popular datasets are GrailQA Gu et al.
+(2021), GraphQA Su et al.
+(2016) and WebQSP Yih et al.
+(2016).
+All of these have the same back-end KB (Freebase).
+For few-shot KBQA transfer, the only available datasets also have only answerable questions Patidar et al.
+(2024).
 GrailQAbility is the only available KBQA dataset with unanswerable questions Patidar et al.
 (2023).
-This was carefully designed starting from GrailQA Gu et al.
-(2021), which has only answerable questions, by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries.
-We use this as one of our targets.
-GraphQA Su et al.
-(2016) is another popular KBQA dataset.
-This has the same back-end KB (Freebase) as GrailQA.
-We create our second target dataset using GraphQA, by replacing its KB with the modified KB in GrailQAbility.
-This introduces unanswerability into GraphQA questions and we label these appropriately as schema-level or data-level unanswerable.
-We name this dataset GraphQAbility.
-We create the two test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from their test sets.
-We create few shots by selecting 100 questions (50 answerable and 50 unanswerable) uniformly at random from the dev set and train set respectively.
+This was constructed starting from GrailQA Gu et al.
+(2021) by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries.
 […]
-WebQSP training set has 2,858 real user questions, which are manually annotated with logical forms.
-This is quite different from GraphQA and GrailQA which contain algorithmically generated logical forms, verbalized by crowd-workers.
-Because of this difference in nature of questions, these are challenging transfer datasets, beyond the unaswerability gap.
-Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
-Schema Level Data Level Schema Level Data Level
+[…]
+WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
+Model Schema Level Data Level Schema Level Data Level
 F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
@@
 For in-domain KBQA with unanswerability, RetinaQA Faldu et al.
-(2024) and Pangu Gu et al.
+(2024) and the unanswerability-adapted version of Pangu Gu et al.
+(2023) are the SoTA models.
+For these, we use the available code.
+[^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in the Appendix (Sec.
+A.8.3).
+FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions.
+KB-Binder Li et al.
+(2023) is the SoTA for in-domain few-shot KBQA.
+Overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al.
 (2024).
-which has been adapted for unanswerable questions, are the two SoTA models.
-For these, we use the available code.
-[^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in Appendix A.10.
-FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions.
-Instead of retrieve-then-generate, KB-Binder Li et al.
-(2023) follows a generate-then-ground approach.
-It is the SoTA for in-domain few-shot KBQA, and overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al.
-(2024).
-We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA since source code is not available.
+We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA.
 To adapt these two baselines for unanswerability, for fair comparison, we modify their logical form generation prompt in the same fashion as PUn for FUn-FuSIC.
@@
 Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, without FUn.
-More details about use of KB-Binder are in the Appendix A.12.
-We use $\mathcal{L}=$gpt-4-0613 for all three models.
-For a fair comparison, we provide all three LLM-equipped models the same aggregated prompt limit for a question.
-To achieve this, to FUn-FuSIC we provide zero-shot generation and $n=4$ FUn iterations, to FuSIC-KBQA-U 5-shot generation, and to KB-Binder 25-shot generation.
+More details about KB-Binder and FuSIC-KBQA are in the Appendix (Sec.
+A.9.1).
+We use $\mathcal{L}=$gpt-4-0613 for all LLM-equipped models.
+For fair comparison, we allocate to all such models the same maximum aggregated prompt length for a question.
+This is satisfied by equipping FUn-FuSIC with zero-shot generation and $n=4$ FUn iterations, FuSIC-KBQA-U with 5-shot generation and KB-Binder with 25-shot generation.
 Though FUn-FuSIC and FuSIC-KBQA allow flexible use of multiple supervised retrievers, for meaningful comparison with RetinaQA, we adapt RetinaQA as retriever for FUn-FuSIC and FuSIC-KBQA-U.
-More details about KB-Binder and FuSIC-KBQA’s generation is in Appendix A.12, and details of FuSIC-KBQA’s retriever along compute infrastructure are in Sec.
-A.9.
+More details about FuSIC-KBQA’s retriever and compute infrastructure are in the Appendix (Sec.
+A.8.2).
 Evaluation Measures:
-Since our primary task is generation of logical forms, our primary focus in evaluation is on logical forms as well.
-For logical form evaluation, the existing exact match measure (EM) Ye et al.
-(2022) works only for s-expression as the language.
-Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression, we propose a new measure that works for both languages.
-This measure, EM-s, considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answers upon execution.
-This is a necessary but not sufficient condition for logical form equivalence.
-So, we compared EM-s with EM in settings where both are applicable, and found $>98\%$ agreement.
-More details are in the Appendix A.3.
-Since logical form evaluation is not completely accurate (logical forms different from the gold one may still be correct), following existing literature Patidar et al.
-(2023); Faldu et al.
-(2024), we also evaluate answers using regular F1 (F1(R)), also lenient F1 (F1(L)), which does not penalize the original answer for the complete KB.
-Answerable Unanswerable
-F1 EM-s F1(L) F1(R) EM-s
-FUn-FuSIC 57.0 46.0 92.0 90.0 62.0
-scUn $\Rightarrow$ sc 63.0 52.0 66.0 64.0 36.0
-w/o syntax 57.0 46.0 92.0 90.0 62.0
-w/o kb-inc 51.7 42.0 92.0 88.0 44.0
-w/o q-lf 47.7 38.0 55.8 48.0 10.0
-w/o ans-inc 55.0 44.0 92.0 90.0 62.0
+For KBQA as semantic parsing task, evaluation of logical forms is primary.
+For this, the existing EM measure Ye et al.
+(2022) is defined only for logical forms represented using s-expressions.
+FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression.
+So we propose a new measure EM-s that checks approximate equivalence for a pair of programs either in SPARQL or s-expression.
+More details are in the Appendix (Sec.
+A.5).
+As in standard KBQA evaluation, we also evaluate answers.
+This is a secondary evaluation for giving the benefit of the doubt for getting the right answer, possibly via a logical form not equivalent to the gold-standard according to EM-s.
+For answer evaluation in KBQA with unanswerability,  Patidar et al.
+(2023) introduced lenient F1, denoted F1(L), in addition to regular F1, denoted F1(R).
+F1(L) relaxes F1(R) by not penalize the original answer for the complete KB.
+Note that obtaining the right answer by chance has much higher probability than for logical forms, particularly for unanswerable questions with NA as the correct answer.
+WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
+Model Answerable Schema L. UnAnswerable Data L. UnAnswerable Answerable Schema Level UnAns Data Level UnAns
+F1 EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
+FUn-FuSIC 74.0 70.0 90.9 87.9 75.8 64.7 64.7 11.8 59.0 48.0 97.1 91.4 80.0 86.7 86.7 26.3
+scUn $\Rightarrow$ sc 74.0 70.0 73.0 72.7 33.3 58.8 52.9 23.5 64.0 54.0 74.4 68.6 22.9 80.0 80.0 13.3
+w/o syntax 67.3 64.0 87.9 90.9 42.4 76.5 70.6 17.7 57.0 46.0 81.0 76.9 26.9 75.8 75.0 16.7
+w/o kb-inc 70.3 66.0 90.9 90.9 9.1 76.5 70.6 29.4 51.7 42.0 100.0 100.0 4.2 79.5 75.0 20.8
+w/o q-lf 71.0 68.0 69.7 63.6 0.0 41.2 35.3 5.9 47.7 38.0 69.5 65.4 07.7 67.0 62.5 20.8
+w/o ans-inc 72.0 70.0 85.9 84.9 33.3 70.6 70.6 23.5 55.0 44.0 80.8 76.9 26.9 79.5 79.2 25.0
 Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on subset of WebQSP $\rightarrow$ GraphQAbility.
@@
 Other rows remove verifiers for syntax error (w/o syntax) (V1), KB inconsistency (w/o kb-inc) (V2), question logical form disagreement (w/o q-lf) (V3) and answer incompatibility (w/o ans-inc) (V4).
+Evaluations are on 100 instances from test sets (50 answerable and 50 unanswerable questions sampled uniformly at random)
 5.2 Unanswerability Setting
 We first address research question R1.
-Performances of the different models for the few-shot transfer setting with unanswerability are recorded in Tab.
+Performances of different models for few-shot transfer with unanswerability are recorded in Tab.
 1.
-Note that the Overall columns determine superiority of one model over another, while the Answerable and Unanswerable columns provide further drill-down for analysis.
 First, we observe that FUn-FuSIC significantly outperforms all baselines in terms of EM-s, and performs at par with FuSIC-KBQA-U and significantly better than all other models in F1.
-However, the other LLM-equipped models do not far surpass supervised models.
-In fact, all 4 models perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility.
-This establishes usefulness of FUn+scUN for few-shot transfer KBQA with unanswerability.
+The other LLM-equipped models are not significantly better than the supervised models.
+All the baselines perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility.
+This establishes usefulness of FUn equipped with scUn for few-shot transfer KBQA with unanswerability, beyond LLM-usage.
 Secondly, each model trades off performance differently between answerable and unanswerable questions.
 RetinaQA, Pangu and also KB-Binder fare better for unanswerable questions, while FUn-FuSIC and FuSIC-KBQA-U fare better for answerable ones.
-However, FUn-FuSIC achieves the best balance between the two.
+However, FUn-FuSIC achieves the best balance across the two subsets.
 We next briefly address research question R2.
@@
 2.
-Models struggle to fare well across both data-level and schema-level unanswerability.
+All models struggle to fare well simultaneously for data-level and schema-level unanswerability.
 FuSIC-KBQA-U performs the best for data-level while performing poorly (in terms of EM-s) for schema-level.
 Conversely, RetinaQA performs well for schema-level, but has poor data-level EM-s.
-FUn-FuSIC performs poorer for data-level compared to schema-level but achieves the best balance by far between the two categories among all models.
+FUn-FuSIC outperforms other models in schema-level unanswerability while being slightly worse in data-level unanswerability.
+But among all models, it achieves the best tradeoff by far across unanswerability categories.
 We next address research question R3.
 Tab.
-3 records performance for different ablations of FUn-FuSIC.
-Note that this experiment is performed on a subset of the test data (50 questions drawn randomly from the each of the answerable and unanswerable categories).
-The biggest benefit, for both answerable and unanswerable questions, comes from Question - Logical Form Disagreement verifier.
-KB Inconsistency and Answer Incompatibility verifiers also make significant contributions to the performance.
-This demonstrates the usefulness of weak verifiers.
-Replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
-Model WebQSP $\rightarrow$ WebQSP $\rightarrow$
-GrailQA-Tech GraphQA-Pop
+3 records the ablation analysis of FUn-FuSIC.
+We see that the KB-Inconsistency verifier (V2) and the Q-LF Disagreement verifier (V3) lead to significant improvements in EM-s.
+The biggest benefit comes from V3 for both answerable and unanswerable questions.
+Without V3, the correct LF for schema-level unanswerability is almost never generated, though answer accuracy stays high, indicating inability to reason.
+Similarly, removing V2 reduces EM-s for schema-level unanswerable questions, with answer accuracy remaining high, indicating that answers are often correct despite flawed logical forms.
+The Answer Incompatibility verifiers (V4) also makes significant contributions to the performance.
+This analysis highlights the necessity of a mix of weak and strong verifiers for structural and semantic validity.
+Beyond verifiers, replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
+WebQSP $\rightarrow$ WebQSP $\rightarrow$
+Model GrailQA-Tech GraphQA-Pop
 FuSIC-KBQA 70.8 52.3
@@
 The models in the top block have prior knowledge of answerability, while those in the bottom block do not.
+[…]
 5.3 Answerable Setting
-We now address research question R4 (answerable-only KBQA).
+We now address research question R4 for answerable-only KBQA transfer.
 We use two datasets from existing literature Patidar et al.
 (2024), including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop).
-[^5] https://github.com/dair-iitd/FuSIC-KBQA/ For consistency with existing literature, here all models use TIARA Shu et al.
-(2022) as the retriever.
-This setting admits two sub-settings: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this prior knowledge.
+[^5] https://github.com/dair-iitd/FuSIC-KBQA/ For enabling comparison with earlier results, we use TIARA Shu et al.
+(2022) as the retriever for all models in this experiment.
+This setting admits two sub-cases: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this knowledge.
 Setting (A) has been studied for KBQA Patidar et al.
 (2024), and FuSIC-KBQA is the established SoTA model, outperforming a host of supervised and LLM-based models adapted for the task.
-In this setting, FUn-FuSIC requires three simple modifications.
-(i) PUn is replaced with prompt for answerability, (ii) In FUn, category of V4b (empty answer) changes from weak verifier to strong verifier, and (iii) scUn is replaced by self-consistency.
+To adapt for this setting, FUn-FuSIC requires three simplifications.
+(i) PUn is replaced with prompt for answerability, (ii) In FUn, V4b (empty answer) is moved from the set of weak verifiers to that of strong verifier, and (iii) scUn is replaced by standard self-consistency.
 The first two rows in Tab.
@@
 FUn-FuSIC significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting.
-This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for KBQA transfer, even without unanswerability.
-In the realistic second setting (B) – not studied before – the models make predictions assuming unanswerability.
+This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for few-shot KBQA transfer, even without unanswerability.
+In the more realistic setting (B), which has not been studied before, the models make predictions assuming unanswerability.
 Here, we evaluate FuSIC-KBQA-U and FUn-FuSIC as in Sec.
-5.2.
+5.2, only there are no truly unanswerable questions.
 The bottom two rows of Tab.
-4 record their performance.
+4 record the performance of the two models in this setting.
 We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin.
@@
 5.4 Error Analysis
-EM-s $<1$ 46.2
+EM-s <1<1< 1 46.2
 Retr. Err. 23.4
 Gen. Err. 22.8
-$l^{*}=$ NK, ${\hat{l}\neq}$ NK 8.4
+$l^{*}=$ NK, $^{\hat{l}\neq}$ NK 8.4
 $l^{*}\neq$ NK, ${\hat{l}=}$ NK 4.6
-$l^{*}\neq$ NK, ${\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
+$l^{*}\neq$ NK, $^{\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
 Table 5: FUn error analysis on WebQSP $\rightarrow$ GraphQAbility.
-$l^{*}$ ${\hat{l}}$ denote gold & generated logical forms.
-Retrieval error means retrieval $r$ is missing $\geq 1$ KB elements (class, relation, entity) necessary for $l^{*}$.
+$l^{*}$ $^{\hat{l}}$ denote gold & generated logical forms.
+Retrieval error means retrieval $r$ is missing 1$\geq$ 1≥ 1 KB elements (class, relation, entity) necessary for $l^{*}$.
 Generation error implies ${\hat{l}}\neq l^{*}$ despite correct retrieval.
-For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions whose logical forms are judged incorrect (EM-s $<1$).
-See results in Tab.
+For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions for which logical forms generated by FUn-FuSIC are incorrect (EM-s <1<1< 1).
+Results are in Tab.
 5.
@@
 (1) Some questions are inherently ambiguous, admitting multiple valid logical forms $l_{1}$ and $l_{2}$ in the original complete KB, though only one is recognized as the gold ($l^{*}=l_{1}$).
-Deletion to introduce unanswerability eliminates $l_{1}$, so we update $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized.
-(2) $l^{*}=l_{1}$, ${\hat{l}}=l_{2}$, where $l_{1}\neq l_{2}$ while being semantically equivalent, but are incorrectly judged non-equivalent by EM-s.
-(3) FUn is genuinely unable to generate $l^{*}$ or any semantic equivalent within its iteration limit.
-We also estimated the accuracy of specific weak verifiers by manual analysis of $\sim$100 instances from test.
-Verifier V3 for Question - Logical Form Disagreement has accuracy of 93.6% overall, with its Query to Natural Language Back-translation component having 90%.
-Verifier V4b for Empty Answer has 67.8% (this is expected since $\sim$25of questions have ($l^{*}\neq$ NK, $A^{*}=$ NA).
+Deletion to introduce unanswerability eliminates $l_{1}$, so that that $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized.
+(2) Here, $l^{*}=l_{1}$ and the prediction ${\hat{l}}=l_{2}$, such that $l_{1}\neq l_{2}$ but are semantically equivalent.
+$l_{1}$ and $l_{2}$ are incorrectly judged non-equivalent by EM-s.
+(3) Here, FUn is unable to generate $l^{*}$ or any semantic equivalent of it within its iteration limit.
 6 Conclusions
-For real-world robust, low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability.
-We have proposed FUn-FuSIC that augments for unanswerability the SoTA few-shot answerable-only KBQA transfer model by (i) iterative repair using feedback from a suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assessing this candidate set to detect unanswerability (and its category) and/or identify the best logical form using self consistency adapted for unanswerability.
-Using two newly created datasets for this novel task, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable few-shot transfer KBQA.
-Error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge and should be a focus of further research.
-7 Limitations
-Since LLM inference involves randomness, ideally experiments should be repeated for multiple runs and results should report averages and error bars.
+For real-world robust and low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability.
+We have introduced a new notion (FUn) of iterative feedback guided repair for answerable as well as unanswerable questions.
+FUn (i) uses feedback from a diverse suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assesses this candidate set to either to detect unanswerability (and its category) or identify the best logical form using self consistency adapted for unanswerability (scUn).
+We propose FUn-FuSIC that replaces the existing the iterative strategy, that assumes answerability of questions, with FUn in the SoTA few-shot answerable-only KBQA transfer model (FuSIC-KBQA).
+Using two newly created datasets for this novel task, we show that FUn-FuSIC significantly outperforms adaptations of FuSIC-KBQA and other SoTA models for this setting, and also for answerable few-shot transfer KBQA.
+Our error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge for KBQA and should be a focus of further research.
+We have made our datasets and other resources public  [^6] https://github.com/dair-iitd/FUn-FuSIC.
+Limitations
+Since LLM inference involves randomness, experiments should ideally be repeated for multiple runs and results should report averages and error bars.
 Unfortunately, we were not able to do this due to the prohibitive cost of GPT-4, and our results are based on single runs.
 While GPT-4 is currently the best performing LLM, it is proprietary as well as expensive.
-Ideally, evaluation should involve open-source freely accessible LLMs as well.
-We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by an open LLM.
-Earlier research has shown that models using Mistral are still able to outperform fully supervised models for answerable few-shot transfer Patidar et al.
+Ideally, evaluation should include open-source freely accessible LLMs as well.
+We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by a less powerful, open LLM.
+Nonetheless, earlier research has shown that models with Mistral instead of GPT-4 still outperform fully supervised models for answerable few-shot transfer Patidar et al.
 (2024).
@@
 That said, following current trends, we expect the ability of open LLMs to steadily improve in the coming years.
-8 Risks
-At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge.
-However, can incorrectly inferring unanswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk?
-While we cannot imagine such a risk at the present time, this may require more thought.
+Risks
+At the highest level, our work reduces risk compared to existing KBQA systems, which when inadequately adapted in a low-resource setting, incorrectly answer unanswerable questions, without acknowledging lack of knowledge.
+However, can incorrectly inferring unanswerability, citing lack of knowledge when knowledge is in fact available, be a new type of risk?
+While we cannot imagine such a risk at the present time, this may require more careful consideration.
 In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
+Acknowledgments
+Mausam is supported by a contract with TCS, grants from IBM, Verisk, Huawei, Wipro, and the Jai Gupta chair fellowship by IIT Delhi.
+Indrajit would like to thank KnowDis AI for supporting his participation in the conference.
+Riya would like to thank Graviton Research Capital for supporting her participation in the conference.
+The authors would like to thank the IIT-D HPC facility for its computational resources.
+We also thank Microsoft Accelerate Foundation Models Research (AFMR) program that provided us access to OpenAI models.
+We are also thankful to Mayur Patidar for helpful discussions.
 Appendix A Appendix
-A.1 KBQA Problem Formulations
-We begin by defining the few-shot transfer learning setting for Knowledge Base Question Answering (KBQA) with unanswerability, and then present our solution for it.
-A.1.1 KBQA: Detailed Background
-A Knowledge Base $G$ consists of a schema with data stored under it.
-The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types.
-The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$.
-We are given a target Knowledge Base $G^{t}$ (consisting of entity types $T^{t}$, relations $R^{t}$, entities $E^{t}$ and facts $F^{t}$) and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns a non-empty answer $A^{t}$ for the question $q^{t}$.
-A.1.2 Few Shot Transfer Learning for KBQA
-In few-shot transfer learning, we are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms in the target domain.
-In addition, we are given access to a related source domain.
-This has a source knowledge base $G^{s}$ consisting of its own types $T^{s}$, relations $R^{s}$, entities $E^{s}$ and facts $F^{s}$, and a much larger source training set $D^{s}$ containing thousands of labeled training examples of questions and corresponding logical forms.
+A.1 KBQA Elaboration
+Here we elaborate on different aspects of the KBQA problem.
+A.1.1 Challenges in Few Shot Transfer Learning for KBQA
 The source and target tasks may differ significantly.
@@
 Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
-A.1.3 KBQA with Unanswerability
-A natural language question q is said to be answerable for a Knowledge Base $G^{t}$ if it has a corresponding logical form $l$ which when executed over $G^{t}$ returns a non-empty answer $A$.
-In contrast, a question $q$ is unanswerable for $G^{t}$, if it either (a) does not have a corresponding logical form that is valid for $G^{t}$, or (b) it has a valid logical form $l$ for $G^{t}$, but which on executing returns an empty answer.
-The first case indicates that $G^{t}$ is missing some schema element necessary for capturing the semantics for $q$.
-The second case indicates that the types $T^{s}$, relations $R^{s}$ is sufficient for $q$, but $G^{t}$ is missing some necessary data elements for answering it.
-In the KBQA with unanswerability task, given a question $q$, if it is answerable, the model needs to output the corresponding logical form $l$.
-and the non-empty answer $A$ entailed by it, and if it is unanswerable, the model either needs to output $NK$ (meaning No Knowledge) for the logical form, or a valid logical form $l$ with $NA$ (meaning No Answer) as the answer.
-Different Types of Unanswerability:
+A.1.2 Different Types of Unanswerability
 Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer.
@@
 However, the (subject, relation, object) path is not connected in the KB.
-A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability
-This setting is a generalization of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$.
-The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form.
-The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
-A.2 Error Checks
+A.2 Algorithm for Self Consistency with Unanswerability
+The high level algorithm for self consistency with Unanswerability (scUn) is described in Algo. 3.
+[…]
+[…]
+[…]
+[…]
+The high level algorithm for assessing confidence in the set of candidate logical forms (assessConf) is described in Algo.
+4.
+We use abbreviations NE and E to indicate non-empty and empty respectively.
+popAnsNE is abbreviation of ”(most) popular answer non-empty”.
+Algorithm 4 $\mbox{assessConf}(q,L,\mathcal{L})$
+[…]
+2:if $(c)$ then
+3:  $l=\mbox{selectBestNE}(q,L^{p},\mathcal{L})$
+4:  $\mbox{{\bf return}}(\mbox{T},l,A^{p})$
+5:end if
+[…]
+7:if $(c)$ then
+8:  $l=\mbox{selectBestE}(q,L^{p},\mathcal{L})$
+9:  $\mbox{{\bf return}}(\mbox{T},l,\texttt{NA})$
+10:end if
+[…]
+A.3 Details of FUn Verifiers
+Here we discuss the verifiers V2, V3 and V4 in more detail.
 (V2) Semantic Error (KB Inconsistency):
@@
 The feedback mentions the type of error and the specifics, e.g., the hallucinated relation, or the incompatible type-relation pair.
-(V3) Question Logical form Disagreement:
+(V3) Question-Logical form Disagreement:
 FUn performs equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline.
@@
 The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$.
-More details about few-shot construction are in Sec.
-A.8.2.
 The feedback mentions lack of equivalence as the type of error.
@@
 As before, the feedback mentions the type of error and the specifics.
-A.3 EM-s: Automated check for equivalence of SPARQL programs
+Knowledge Base Logical Forms NL Questions
+[…]
+SrcAv/TgtAv Cosine Sim SrcAv/TgtAv
+[…]
+[…]
+Table 6: Statistics for different source and target (test set) KBQA task pairs in terms of the knowledge base, logical forms and natural language questions.
+‘Domain(JS)’ is Jensen Shannon (JS) divergence between domain distribution of questions, ‘New Rel%’ is percentage of questions in target with new (unseen) relations, ‘Function (JS)’ is JS-divergence between distributions over functions in logical forms, ‘#\##Relations’ shows the source average and the target average for number of relations per logical form, ‘Src-Tgt Cosine Sim’ is average minimum cosine distance between source and target questions and ‘#\##Tokens’ shows the source average and target average of number of tokens per question.
+A.4 Additional Dataset Statistics
+[…]
+Dataset Schema level Data level
+UnAns UnAns
+[…]
+[…]
+Table 7: Percentages of schema and data level unanswerability among unanswerable questions.
+A.5 EM-s: Automated Approximate Equivalence Check for SPARQL
 As has been observed in Patidar et al.
@@
 However, FUn-FuSIC generates SPARQL queries instead.
-Directly comparing program equivalence between two sparql queries is an undecideable problem  [^6] https://users.dcc.uchile.cl/ cgutierr/papers/expPowSPARQL.pdf.
+Directly comparing program equivalence between two SPARQL queries is an undecidable problem  [^7] https://users.dcc.uchile.cl/~cgutierr/papers/expPowSPARQL.pdf.
 Patidar et al.
-(2024) suggests a semi-automatic strategy for comparison of sparql queries.
+(2024) suggest a semi-automatic strategy for comparison of sparql queries.
 We propose a completely automatic metric for SPARQL equivalence check.
-Two sparql queries are equivalent by the EM-s check if (a) the relations occuring in the two queries are same.
-(b) the entities occuring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same.
+Two SPARQL queries are equivalent by the EM-s check if (a) the relations occurring in the two queries are same.
+(b) the entities occurring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same.
 Note that the EM-s check is necessary, but not sufficient for two SPARQL queries to be equivalent.
-Since these are a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable and found $>98\%$ agreement.
-A.4 Performance across different categories of unanswerability
-We dive deeper into the different categories of unanswerability, as explained in Patidar et al.
+[…]
+A.6 Model Evaluation: Additional Details
+In this section, we do a deeper evaluation of performance of different models across different categories of unanswerability, as explained in Patidar et al.
 (2023).
-There exist 2 broad categories of unanswerability- schema level unanswerability(absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability(absence of facts or intermediate entities of the logical form path on the KB).
-We expect that (a) due to the poor ability of supervised models to generalize in transfer learning settings, RetinaQA will be struggle to generate correct logical forms for data level unanswerable questions (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions.
-However, since it is biased to return incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions.
+There are two broad categories of unanswerability — schema level unanswerability (absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability (absence of facts or intermediate entities of the logical form path on the KB).
+We expect that (a) due to poor ability of supervised models to generalize in transfer learning settings, RetinaQA will struggle to generate correct logical forms for data level unanswerable questions, and (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions.
+However, since it is biased towards returning incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions.
 (c) FUn-FuSIC should be able to maintain the performance of FuSIC-KBQA on data level unanswerable questions to a large extent, while significantly improving the performance on schema level unanswerable questions.
-Based on performance on the WebQSP $\rightarrow$ GrailQAbility and WebQSP $\rightarrow$ GraphQAbility datasets, we validate the that the trends are indeed as expected.
-A.5 Cost Analysis
-FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as U-FuSIC and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16.
+[…]
+A.7 FUn Cost Analysis for Proprietary LLMs
+FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as FuSIC-KBQA-U and FUn-FuSIC rerank the classes, relations and paths.
+The total cost for reranking for one question is \$0.16.
 The cost for generation of logical form from a prompt with 5 in-context examples is \$0.16.
-Thus, the approximate cost for inference of one question by U-FuSIC is \$0.32.
+Thus, the approximate cost for inference of one question by FuSIC-KBQA-U is \$0.32.
 The cost for generation of logical form from a prompt with 0 in-context examples is \$0.04.
 The cost of checking whether two natural language questions are equivalent or not, using few-shot exemplars and chain of thought prompting is also \$0.04.
-The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48. The average cost over 50 randomly sampled questions from the test set is around \$0.34.
+The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48.
+The average cost over 50 randomly sampled questions from the test set is around \$0.34.
 Hence, the two models are comparable in terms of cost.
-A.6 FuSIC-KBQA Details
-Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA.
+A.8 Model Adaptation Details
+Here we discuss adaptation details for the models that we have built upon (FuSIC-KBQA), used as retrievers (RetinaQA) and for comparison.
+A.8.1 FuSIC-KBQA Details
+Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA Patidar et al.
+(2024).
 FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the SPARQL query.
-A.7 FUn-FuSIC Prompts
+Since no code is available for this model, we use our own implementation based on the description in the paper. For FuSIC-KBQA, and FUn-FuSIC we use LLM $temperature=0$.
+A.8.2 Training Details for Supervised Models
+We use Hugging Face  Wolf et al.
+(2020), PyTorch Paszke et al.
+(2019) for our experiments and use the Freebase setup specified on github  [^8] https://github.com/dki-lab/Freebase-Setup .
+We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM.
+For training the discriminator module of RetinaQA, we require 2 GPUs.
+(1) For the answerable experiments, we use the supervised models as specified in Patidar et al.
+(2024).
+(2) For the unanswerability experiments, we train all models from scratch.
+(a) We use RnG-KBQA entity linker  [^9] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments.
+(b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^10] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License).
+The number of training epochs is determined by the performance of the model over the answerable questions in the dev set.
+(c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^11] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^12] https://github.com/dair-iitd/RetinaQA.
+A.8.3 Inference Details for Supervised Models
+We train all RetinaQA components on the source WebQSP’s training set, using the corresponding target domain’s dev set as a validation set for early stopping.
+In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability.
+We again use the target dev sets for this.
+We use the dev set in RetinaQA, during discriminator inference for different purposes.
+(A) Determining how to best utilize the candidate paths.
+The possibilities are (i) not providing candidate paths, (ii) providing candidate paths in GrailQA format, and (iii) providing candidate paths in WebQSP format.
+We select the best alternative based upon the performance of the model over the dev set.
+For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best.
+(B) Determining the threshold value.
+RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK.
+We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
+A.9 Pangu Adaptation Details
+Similar to RetinaQA, we train all Pangu components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping.
+We use one GPU for training.
+Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability.
+Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK.
+We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
+A.9.1 KB-Binder Adaptation Details
+For KB-Binder, we make use of publicly available code  [^13] https://github.com/ltl3A87/KB-BINDER (MIT License).
+We use self-consistency and majority voting with 6 examples, as in the experiments in the paper.
+In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data.
+We restrict its retrieval to our target training set $D_{t}$ with 25 examples.
+KB-Binder reports experiments using code-davinci-002 as the LLM.
+For consistency and fair comparison, we replace this with gpt-4-0613 as in other LLM-equipped models.
+KB-Binder generates logical forms in s-expression, which we preserve.
+WebQSP $\rightarrow$ WebQSP $\rightarrow$
+Model GrailQAbility GraphQAbility
+V3(a) 92 96
+V3(b) 94 90
+V3(c) 92 94
+V3(overall) 88 90
+V4(b) 75 68
+Table 8: Accuracy of weak verifier on the two datasets by manual analysis of 100 instances from the test sets.
+A.10 Accuracy of Weak Verifiers
+Accuracy of the two weak verifiers are recorded in detail in Tab. 8.
+A.11 FUn-FuSIC Prompts
 Here we provide details of various prompts used by FUn-FuSIC.
-A.8 PUn prompt
-The first prompt is for Prompting for Unanswerability (PUn).
+A.11.1 PUn prompt
+The following prompt is for Prompting for Unanswerability (PUn).
 Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively.
@@
 sparql:
-A.8.1 FUN prompt
+A.11.2 FUN Prompts
+The following is the prompt used by FUn for accommodating feedback from verifier V1.
 Syntax error(V1) Feedback Correct the syntax of the following sparql query.
@@
 Virtuoso error: word AND not defined
+The following is the prompt used by FUn for accommodating feedback from verifier V2.
 KB Inconsistency(V2) Feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query.
@@
 DO NOT APOLOGIZE - just return the best you can try.
+The following is the prompt used by FUn for accommodating feedback from verifier V3.
 Question Logical form disagreement(V3) feedback The question that you answer is NOT same as what you’ve been asked for!
@@
 DO NOT APOLOGIZE - just return the best you can try.
-The next three prompts fall under the answer incompatibility feedback
+The following three prompts are used by FUn for accommodating feedback from verifier V4.
 Answer Inconsistency(V4b) feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
@@
 Please reconstruct the query using same context and constraints.
-A.8.2 Prompt for Verifier V3 Question Logical Form Agreement
-The few shots provided for verifying question logical form agreement come from the few shots provided in the target domain $D^{t}$.
+A.11.3 Prompt for Question Logical Form Agreement Verifier (V3)
+The few shots provided for verifying question logical form agreement are derived from $D^{t}$.
 We obtain positive samples from the dataset $D^{t}$ directly, using the questions and gold logical forms.
-For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over the dataset $D^{t}$.
+For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over $D^{t}$.
 Then we consider those questions for which the predicted logical form is different from the gold logical form.
-Firstly, we perform back-translation to obtain natural language question from the logical form
+First, we perform back-translation to obtain natural language question from the logical form using the following prompt.
 Naturalization of variable names(V3(i)) change the sparql query to have variable names representative of what objects they refer to.
@@
 […]
-We use few-shot LLM prompting to obtain the explanation for why the question and logical form agree or disagree.
-These few-shots, for obtaining the explanation are dataset independent, and are manually written.
+We use few-shot LLM prompting to obtain the explanation for question and logical form agreement or disagreement.
 Explanation Generation Prompt Explain why the two questions are different.
@@
 explanation:
-Some examples of few shots are provided below-
-Question Logical Form Sgreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals?
+These few-shots for obtaining the explanation are dataset independent, and are manually written. Some examples of few shots are below.
+Question Logical Form Agreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals?
 Question originally asked: name the cricket player who has 31 odi stumps.
@@
 explanation:
-A.8.3 Prompt in scUn for Logical Form Selection
+A.11.4 scUn Prompt for Logical Form Selection
 We store the back-translated natural language questions along with each predicted logical form.
-Here, we prompt the LLM to select the closest back-translated natural language question to the original question.
+We use the following prompt to select the closest back-translated natural language question to the original question.
 Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director?
@@
 Please explain your answer as well
-A.9 Supervised Models training details
-We use Hugging Face  Wolf et al.
-(2020), PyTorch Paszke et al.
-(2019) for our experiments and use the Freebase setup specified on github 777 https://github.com/dki-lab/Freebase-Setup .
-We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM.
-For training the discriminator module of RetinaQA, we require 2 GPUs.
-(1) For the answerable experiments, we use the supervised models as specified in Patidar et al.
-(2024).
-(2) For the unanswerability experiments, we train all models from scratch.
-(a) We use RnG-KBQA entity linker  [^8] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments.
-(b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^9] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License).
-The number of training epochs is determined by the performance of the model over the answerable questions in the dev set.
-(c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^10] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^11] https://github.com/dair-iitd/RetinaQA.
-A.10 Supervised Models inference details
-We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping.
-In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability.
-We again use the target dev sets for this.
-We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths.
-We might (i) not provide candidate paths (ii) provide candidate paths in GrailQA format (iii) provide candidate paths in WebQSP format.
-We select the best alternative based upon the performance of the model over the dev set.
-For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best.
-(b) determining the threshold value.
-RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK.
-We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
-A.11 Pangu Adaptation Details
-Similar to RetinaQA, We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping.
-We use 1 GPU for training.
-Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability.
-Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK.
-We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
-A.12 KB-Binder and FuSIC-KBQA Adaptation Details
-For KB-Binder, we make use of publicly available code  [^12] https://github.com/ltl3A87/KB-BINDER (MIT License).
-We use self-consistency and majority voting with 6 examples, as in the experiments in the paper.
-In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data.
-We restrict its retrieval to our target training set $D_{t}$ with 25 examples.
-KB-Binder reports experiments using code-davinci-002 as the LLM.
-For consistency and fair comparison, we replace these with gpt-4-0613 as in our approach.
-This model generates logical forms in s-expression, which we preserve.
-For FuSIC-KBQA Patidar et al.
-(2024), since no code is available for this model, we use our own implementation based on the description in the paper.
-For FuSIC-KBQA, and FUn-FuSIC we use $temperature=0$.
-A.13 Real Example of FUn
-Question: Where all were the Olympics held before 2024?
-Logical Form 1 ?y{ ?x olympics.venue ?y
-    ?y olympics.year ?z
-    FILTER(?z = 2024)}
-Feedback 1 Type: Logical Form Semantic Error Description: ?y is assigned types [’venue’, ’olympics’] by relations [’olympics.venue’, ’olympics.year’]. These types are mutually incompatible.
-Logical Form 2 ?y{ ?x olympics.venue ?y
-    ?x olympics.year ?z
-    FILTER(?z = 2024)}
-Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked returns all places where the Olympics were held before 2024.
-However, the question we answer returns the places where the Olympics were held in 2024.
-Logical Form 3 ?y{ ?x olympics.venue ?y
-    ?x olympics.year ?z
-    FILTER(?z < 2024)}
+A.12 Example of FUn Iterations
+The following shows 3 Fun iterations for an example question from the test set of WebQSP $\rightarrow$ GrailQAbility dataset.
+Question: what is the musical genre of the recording who m i (feat. 일리닛, new champ, myk)?
+Entities: who m i (feat. 일리닛, new champ, myk) m.0123lk0s
+Logical Form 1 ?x {
+    :m.0123lk0s :music.genre.recordings ?x
+    ?x :type.object.type :music.genre}
+Feedback 1 Type: Logical Form Semantic Error Description: The types of relations don’t match for entity in the query.
+The assigned relation types by [’music.genre.recordings’] are [’music.genre’].
+These types are not associated with this entity in the KB.
+Logical Form 2 ?y {
+    :m.0123lk0s :music.recording.artist ?x
+    ?x :music.artist.genre ?y
+    ?y :type.object.type :music.genre}
+Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked genre of the song. However, the question we answer returns genre associated with artist of the song.
+Logical Form 3 ?x { ?x
+    :music.genre.recordings :m.0123lk0s
+    ?x :type.object.type :music.genre}
 Feedback 3 Status: Passes All Checks Details: lf3 satisfies syntax, logical form semantic, natural language semantic and execution checks.
-A.14 Algorithm for Self Consistency with Unanswerability
-The high level algorithm for self consistency with Unanswerability (scUn) is described in Algo. 3.
-Algorithm 3 ScUn$(q,L,t,\mathcal{L})$
-1:$(c,l,A)=\mbox{assessConf}(q,L,t,\mathcal{L})$
-2:if ($\mbox{{\bf if} }(c)$, $\mbox{{\bf return}}(l,A)$
-[…]
-The high level algorithm for assessing confidence in the set of candidate logical forms (assessConf) is described in Algo. 4.
-Algorithm 4 $\mbox{assessConf}(q,L,t,\mathcal{L})$
-1:$(c,L^{p},A^{p})=\mbox{popAnsNE}(L,t)$
-2:if $(c)$ then
-3:     $l=\mbox{selectBestNE}(q,L^{p},\mathcal{L})$
-4:     $\mbox{{\bf return}}(\mbox{True},l,A^{p})$
-5:end if
-6:$(c,L^{p})=\mbox{popAnsE}(L,t)$
-7:if $(c)$ then
-8:     $l=\mbox{selectBestE}(q,L^{p},\mathcal{L})$
-9:     $\mbox{{\bf return}}(\mbox{True},l,\texttt{NA})$
-10:end if
-[…]